### v1

# CKD: Contrastive Knowledge Distillation from A Sample-wise Perspective

Abstract
In this paper, we present a simple yet effective contrastive knowledge distillation approach, which can be formulated as a sample-wise alignment problem with intra- and inter-sample constraints. Unlike traditional knowledge distillation methods that concentrate on maximizing feature similarities or preserving class-wise semantic correlations between teacher and student features, our method attempts to recover the “dark knowledge” by aligning sample-wise teacher and student logits. Specifically, our method first minimizes logit differences within the same sample by considering their numerical values, thus preserving intra-sample similarities. Next, we bridge semantic disparities by leveraging dissimilarities across different samples. Note that constraints on intra-sample similarities and inter-sample dissimilarities can be efficiently and effectively reformulated into a contrastive learning framework with newly designed positive and negative pairs. The positive pair consists of the teacher’s and student’s logits derived from an identical sample, while the negative pairs are formed by using logits from different samples. With this formulation, our method benefits from the simplicity and efficiency of contrastive learning through the optimization of InfoNCE, yielding a run-time complexity that is far less than $\mathcal{O}(n^{2})$, where $n$ represents the total number of training samples. Furthermore, our method can eliminate the need for hyperparameter tuning, particularly related to temperature parameters and large batch sizes. We conduct comprehensive experiments on three datasets including CIFAR-100, ImageNet-1K, and MS COCO. Experimental results clearly confirm the effectiveness of the proposed method on both image classification and object detection tasks. Our source codes will be publicly available at https://github.com/wencheng-zhu/CKD.
Index Terms:
Knowledge distillation, contrastive learning, intra- and inter-sample constraint, sample-wise alignment
I Introduction
Knowledge distillation (KD) [68, 15, 14] compresses knowledge without significant performance degradation, making it possible to deploy resource-intensive models on lightweight devices [40, 74, 19]. Given the practical benefits, knowledge distillation has garnered much attention over the past decade, and numerous knowledge distillation approaches have been proposed and further introduced into a wide spectrum of tasks [10, 5, 75], such as image classification [39, 16, 71], object detection [87, 82, 61], and segmentation [45, 62, 81]. While considerable improvements have been achieved, existing knowledge distillation methods still struggle to determine what knowledge to learn and how to learn it [66, 2, 53]. Here, we posit that the student model can effectively replicate the output of the teacher model when each sample is appropriately aligned.
Figure 1: Illustration of three typical approaches, including classic KD [25], CRD [70], and our CKD. Classic KD emphasizes feature similarities between teacher and student features, while CRD focuses on class-wise semantic information. In contrast, the proposed CKD conducts similarity alignment and captures sample-wise semantic structures simultaneously.
Hinton et al. [25] propose a seminal work on knowledge distillation, with the goal of reducing the Kullback-Leibler (KL) divergence between teacher and student logits. This influential research has led to a surge in related studies, most of which can be categorized into logit-based and feature-based approaches. Logit-based approaches distill the output logits of teachers [83, 35]. Representative methods in this category are vanilla KD [25], DML [89], DIST [26], DKD [90], etc. Following vanilla KD [25], there has been a lack of research on logit distillation, mainly because of its inferior performance relative to feature distillation. However, logits contain compact information that is easier to distill. Furthermore, logit-based methods can take advantage of low computational and storage requirements [34, 37, 86]. Our approach also employs logit distillation.
Feature-based approaches seek to replicate features extracted from intermediate layers of a teacher model [42, 20]. Typical methods include FitNet [59], OFD [24], ReviewKD [7], etc. These methods have advanced the field of knowledge distillation. However, due to rote memorization, they often suffer from severe overfitting that leads to low training losses but high test losses. Given that, RKD [55] resembles high-order structural relations among features instead of feature similarities. SP [72] preserves pairwise similarities in representation space. Although feature-based approaches effectively capture feature similarity structure, they may somewhat overlook semantic knowledge. CRD [70] innovatively distills structural representations through contrastive learning. To handle the requirement for large batch sizes in contrastive learning, CRD implements a memory buffer to enable the sampling of many negative samples, resulting in less-than-ideal training efficiency, particularly for dense prediction tasks.
To address the aforementioned structure-preserving issues, we present a Contrastive Knowledge Distillation approach, dubbed CKD, which treats knowledge distillation as an intra- and inter-sample alignment problem. Fig. 1 briefly illustrates the main idea of the proposed approach in comparison with two representative approaches. We clearly observe that the proposed approach pulls logits from the same sample together while pushing logits from different samples apart. Specifically, our method assumes that a student model can well mimic a teacher model by distilling sample-wise information. This idea is straightforward and reasonable. On the one hand, previous works [59, 7] have shown tremendous success by maximizing intra-sample feature similarities. Inspired by these methods, we directly minimize logit differences within the same sample. On the other hand, solely relying on an intra-sample similarity constraint can easily lead to overfitting [57, 77]. To this end, we dive into the inter-sample constraints among different samples for additional information. Finally, we integrate these two constraints into the loss function.
Formally, the intra- and inter-sample constraints can be cast as a sample-wise contrastive learning problem. Compared to existing contrastive learning approaches [70, 4] that preserve the class-wise semantic structure, the sample-wise contrastive framework can distill teacher knowledge in an unsupervised way without label information. Moreover, our method precisely aligns teacher and student logits by exploiting both similarities and dissimilarities within each sample rather than each class. From the sample-wise perspective, this leads to a relatively small number of associated negative samples, without resorting to large batch sizes for sampling the whole negative space. Thus, the sample-wise formulation enables our approach to maintain low computational complexity in the absence of class-wise information. Furthermore, our method employs logit distillation over feature distillation for two main reasons. One reason is that logits contain high-level and compact semantic information, which is particularly suitable for contrastive learning. The other reason is that the dimension of logits, crucial for negative sampling, is lower. Notably, our method is model-agnostic and can be applied to feature distillation methods.
To sum up, the contributions of this paper are summarized into three aspects:
  1. 1.
    We present a contrastive knowledge distillation approach from the sample-wise perspective. Our method simultaneously conducts intra-sample similarity and inter-sample dissimilarity alignment between teacher and student logits.
  2. 2.
    We formulate these two intra-sample and inter-sample constraints as a contrastive learning problem with well-designed positive and negative pairs. By using the proposed contrastive formulation, our method can be efficiently and effectively trained.
  3. 3.
    We evaluate the proposed approach on benchmarks for image classification and object detection. Without bells and whistles, our approach improves the accuracy of vanilla KD by more than 1.4% and 0.98% both on the CIFAR-100 and ImageNet-1K datasets.
II Related Work
In this section, we first outline recent progress in knowledge distillation [49, 41, 38]. Then, we review related contrastive learning approaches.
II-A Knowledge Distillation
Knowledge distillation is a representative model compression technique [65, 49, 80] that enables small models to attain strong performance of large models [1, 52, 48]. As a pioneer, Hinton et al. [25] introduced the concept of knowledge distillation that transfers “dark knowledge” by minimizing the KL divergence between teacher logits and student logits. Over the past few years, a rich line of knowledge distillation methods has been proposed [30, 11, 12]. These knowledge distillation methods come in two flavors: logit distillation and feature distillation [50, 33].
II-A1 Logit Distillation
Earlier logit distillation approaches attempt to enhance generalization performance through regularization and optimization techniques. However, there has been limited research on early-stage logit distillation due to inferior performance. DML [89] improved the generalization abilities of networks by collaboratively training an ensemble of models through mutual learning. Instead, TAKD [51] employed multi-step distillation via intermediate-sized networks, effectively bridging the knowledge transfer gap between teacher and student networks. ICKD-C [44] utilized grid-level inter-channel correlation to capture the intrinsic distribution and diversity of feature representations. GLD [29] employed a local spatial pooling layer for the simultaneous extraction of finely localized knowledge and holistic representations. DIST [26] leveraged both inter-class and intra-class relations to align probabilistic distributions and reinforce semantic similarities, respectively. DKD [90] decoupled logits into a target class and all non-target classes, thereby enhancing both the effectiveness and flexibility of knowledge transfer. Auto-KD [35] firstly introduced an automated search strategy by employing Monte Carlo tree search to optimize knowledge distillation architectures. CTKD [39] modulated the task difficulty levels of a student network by using a dynamic and learnable temperature setting. MLKD [27] also proposed a multi-level logit distillation framework, which distills instance-level, batch-level, and class-level information. LSKD [67] considered temperature as the standard deviation of logit and applied logit standardization before the softmax prediction. Besides, there are several other studies [57, 9] devoted to explaining the basic principle of knowledge distillation. Our method belongs to logit distillation that leverages contrastive information between the logits of the teacher and student models.
Figure 2: The overall architecture of the proposed CKD. The newly designed triplets are introduced for similarity and semantic alignment. For example, we classify all other samples as negative, regardless of their class. In this formulation, training efficiency can be improved by reusing negative samples.
II-A2 Feature Distillation
Feature distillation emphasizes the maximization of similarities between intermediate representations. FitNet [59] proposed a feature-based approach that extracted trainable student representations to predict teacher representations. But, the excessive proximity between samples often results in overfitting. Hence, AT [31] mimicked attention maps of a teacher model to enhance student performance. Instead, RKD [55] transferred the relational structure among data examples by incorporating both distance-wise and angle-wise distillation objectives. IRG [46] distilled not just instance features but also the often-overlooked instance relationships and cross-layer feature transformations. CCKD [56] captured both instance-level information and correlations between multiple instances for knowledge transfer. SP [72] introduced a similarity-preserving loss to maintain the activation pattern similarities of input pairs. OFD [24] combined a new margin ReLU, an optimized distillation feature position, and a selective L2 distance to enhance network compression. CRD [70] was the first contrastive distillation approach that preserves structural representations among features. ReviewKD [7] distilled feature representations from cross-level connection paths to guide the learning of a student model. SimKD [6] reused the teacher’s classifier and introduced an additional projector to align student features. CAT-KD [17] enhances knowledge distillation’s interpretability by distilling class activation maps. This approach empowers the student model to identify discriminative regions, improving its performance. Feature-based approaches usually outperform logit-based approaches at the expense of additional computational and storage overhead.
II-B Contrastive Learning
Contrastive learning stands out as a powerful and popular branch of self-supervised learning, and plenty of efforts have been devoted to contrastive learning [3, 23]. We merely review relevant works concerned with NCE and InfoNCE.
Gutmann et al. [18] firstly proposed the theory of noise-contrastive estimation, in which the model learns to discriminate between samples from the actual data distribution and those from a noise distribution, thereby revealing distinct features of the data. In noise-contrastive estimation, increasing the volume of negative samples can largely reduce reliance on the quality of the noise distribution. Sohn et al. [64] became the first to introduce the multi-class N-pair loss for deep metric learning. This approach eliminates the necessity to recalculate embedding representations in every iteration by allowing the reuse of embedding representations from negative samples. InstDisc [78] incorporated a memory bank mechanism that retains instance features from the preceding update step, facilitating negative sampling through integration with this non-parametric memory bank. CPC [54] popularized the InfoNCE loss for contrastive representation learning, which enhances similarities between a sample and its corresponding positive samples while decreasing similarities with negative samples. Likewise, CMC [69] leveraged a contrastive loss function to optimize the mutual information across different view representations of the same sample. MoCo [21] implemented a moving average of model parameters to update a dynamic dictionary using a queue. SimCLR [8] employed data augmentation to generate varied inputs of the same image and employed a contrastive loss to enhance the similarity between representations with different augmentations. CLIP [58] proposed a large-scale multimodal model pre-trained through the joint training of a text encoder and an image encoder with the InfoNCE loss. Previous methods have witnessed the success of contrastive learning. However, most contrastive learning approaches require substantial batch sizes of negative samples to optimize training. In this paper, we harness the potential of contrastive learning and design sample-wise triples for efficient and effective training.
III Approach
In this section, we first describe some notations. Then, we delve into contrastive knowledge distillation. Lastly, we provide a discussion about the proposed method.
III-A Notations
Knowledge distillation tasks aim to transfer knowledge from the teacher model $\mathcal{T}$ to the student model $\mathcal{S}$. To be specific, given inputs $\mathcal{X}=[\boldsymbol{x}_{0},\boldsymbol{x}_{1},\ldots,\boldsymbol{x}_{n-1}]$, we calculate teacher logits $\boldsymbol{T}=\mathcal{T}\left(\mathcal{X};\theta_{\mathcal{T}}\right)$ and student logits $\boldsymbol{S}=\mathcal{S}\left(\mathcal{X};\theta_{\mathcal{S}}\right)$, where $\boldsymbol{T}=[\boldsymbol{t}_{0},\boldsymbol{t}_{1},\ldots,\boldsymbol{t}_{n-1}]\in\mathbb{R}^{c\times n}$ and $\boldsymbol{S}=[\boldsymbol{s}_{0},\boldsymbol{s}_{1},\ldots,\boldsymbol{s}_{n-1}]\in\mathbb{R}^{c\times n}$. Here, $\theta_{\mathcal{T}}$ and $\theta_{\mathcal{S}}$ are model parameters. $\mathbf{t}_{i}$ and $\mathbf{s}_{i}$ represent the teacher and student logits for the $i$-th sample in which $i\in\mathcal{D}$, $\mathcal{D}=[1,2,\ldots,n-1]$. $n$ denotes the sample count and $c$ signifies the logit dimension that corresponds to the number of classes. Formally, the objective of knowledge distillation approaches can be written as:
$\mathcal{L}=\mathcal{L}_{\text{Task}}+\alpha\mathcal{L}_{\text{KD}}.$ (1)
$\mathcal{L}_{\text{Task}}$ represents the standard cross-entropy loss for image classification, and $\alpha$ balances the classification loss $\mathcal{L}_{\text{Task}}$ and the distillation loss $\mathcal{L}_{\text{KD}}$. In what follows, we elaborate on the proposed knowledge distillation loss $\mathcal{L}_{\text{KD}}$.
III-B Contrastive Knowledge Distillation
Generally, traditional knowledge distillation methods highlight class-level contrastive information and capture intra-class and inter-class relationships. Differently, our method focuses on sample-level representations. To be specific, the proposed approach aligns a student logit $\mathbf{s}_{i}$ with its corresponding teacher logit $\mathbf{t}_{i}$ while pushing it apart from other negative logits $\mathbf{s}_{j}$. To achieve this goal, we endeavor to distill not only intra-sample numerical similarities but also inter-sample semantic structure. Fig. 2 presents the overall architecture of CKD. We can see that our method conducts intra-sample distillation and inter-sample distillation. Afterward, we describe these constraints in detail.
III-B1 Intra-Sample Distillation
Class-wise methods usually minimize intra-class distances among samples for knowledge transfer. By adhering to these methods, the proposed approach maximizes similarities between teacher and student logits for each sample while leveraging sample-wise information. Here, we utilize the intra-sample loss, denoted as $\mathcal{L}_{\text{intra}}$, to optimize parameters of the student model $\mathcal{S}$, where $d(\cdot,\cdot)$ represents a distance metric. Hence, we derive:
$\mathcal{L}_{\text{intra}}$ $=\frac{1}{n}\sum_{i=0}^{n}d\left(\boldsymbol{t}_{i},\boldsymbol{s}_{i}\right)$ (2)
$=\underset{\boldsymbol{x}_{i}\sim\mathcal{X}}{\mathbb{E}}{d\left(\boldsymbol{t}_{i},\boldsymbol{s}_{i}\right)}.$
For $\boldsymbol{x}_{i}\in\mathcal{X}$, we obtain $\boldsymbol{t}_{i}=\mathcal{T}\left(\boldsymbol{x}_{i};\theta_{\mathcal{T}}\right)$ and $\boldsymbol{s}_{i}=\mathcal{S}\left(\boldsymbol{x}_{i};\theta_{\mathcal{S}}\right)$, where $\boldsymbol{x}_{i}\sim\mathcal{X}$ denotes a data point $\boldsymbol{x}_{i}$ sampled from $\mathcal{X}$. The loss function $\mathcal{L}_{\text{intra}}$, defined in Eq. 2, forces the student logit $\boldsymbol{s}_{i}$ to approximate its analytical solution $\boldsymbol{t}_{i}$. Previous studies [59, 55] have shown that continuously minimizing the gap between teacher and student representations can lead to overfitting and reduced generalization to unseen data. This is evident in the significant discrepancy between training and validation losses, especially when relying solely on intra-sample similarities.
(a) $d(\mathbf{t}_{i},\mathbf{s}_{i})<d(\mathbf{t}_{j},\mathbf{s}_{i})$
(b) $d(\mathbf{t}_{i},\mathbf{s}_{i})>d(\mathbf{t}_{j},\mathbf{s}_{i})$
Figure 3: Visualization of the intra-sample alignment. For the sake of simplicity, we assume that the dimension of logits is 3. The spherical surface represents $\mathbf{s}_{i}$ satisfying the condition $\epsilon_{i}=\mathbf{t}_{i}-\mathbf{s}_{i}$ for the fixed $\epsilon_{i}$.
(a) $\left(\boldsymbol{s}_{i},\boldsymbol{t}_{i},\boldsymbol{s}_{j}\right)$
(b) $\left(\boldsymbol{s}_{i},\boldsymbol{t}_{i},\boldsymbol{t}_{j}\right)$
(c) $\left(\boldsymbol{s}_{i},\boldsymbol{t}_{i},\boldsymbol{t}_{j}\right)$
Figure 4: Visualization of two categories of triples. (a) depicts $\left(\boldsymbol{s}_{i},\boldsymbol{t}_{i},\boldsymbol{s}_{j}\right)$, while (b) and (c) show $\left(\boldsymbol{s}_{i},\boldsymbol{t}_{i},\boldsymbol{t}_{j}\right)$. The blue line indicates that $\boldsymbol{s}_{i}$ approaches $\boldsymbol{t}_{i}$, and the red line represents that $\boldsymbol{s}_{i}$ is far from $\boldsymbol{s}_{j}$ or $\boldsymbol{t}_{j}$.
Fig. 3 provides a detailed analysis and visualization of the intra-sample constraint on logits $\boldsymbol{t}_{i}$, $\boldsymbol{t}_{j}$, and $\boldsymbol{s}_{i}$. The error between the $i$-th teacher and student logits is represented as $\epsilon_{i}=\boldsymbol{t}_{i}-\boldsymbol{s}_{i}$, where $\boldsymbol{t}_{i}\neq\boldsymbol{t}_{j}$. The objective $\mathcal{L}_{\text{intra}}$ cannot guarantee $\forall I,d(\boldsymbol{t}_{i},\boldsymbol{s}_{i})>d(\boldsymbol{t}_{j},\boldsymbol{s}_{i})$. Therefore, there exist errors $\epsilon_{i}$, such that ${\exists}i,d(\boldsymbol{t}_{i},\boldsymbol{s}_{i})>d(\boldsymbol{t}_{j},\boldsymbol{s}_{i})$. This inequality suggests that the student logit $\boldsymbol{s}_{i}$ may be closer to a teacher logit $\boldsymbol{t}_{j}$ than its corresponding teacher logit $\boldsymbol{t}_{i}$.
Formally, we assume that there are hyperplanes $\mathcal{H}^{(k)},k\in\mathbb{N}$ related to $\boldsymbol{s}_{i}^{(k)}\in\mathcal{H}^{(k)},i\in\mathcal{D}$ with the same loss value, that is:
$\forall k,h\;\;\sum_{i\in\mathcal{D}}d\left(\boldsymbol{t}_{i},\boldsymbol{s}_{i}^{(k)}\right)=\sum_{i\in\mathcal{D}}d\left(\boldsymbol{t}_{i},\boldsymbol{s}_{i}^{(h)}\right).$ (3)
The hyperplane $\mathcal{H}^{(k)}$ is defined as:
$\mathcal{H}^{(k)}:\mathbf{w}^{(k)}\boldsymbol{s}^{(k)}_{i}+b^{(k)}=0,\;i\in\mathcal{D},$ (4)
where $\boldsymbol{w}\in\mathbb{R}^{c}$ is a normal vector and $b$ is a bias.
Consider a two-dimensional plane $y=x$ and a teacher logit $\left[0.5,0.5\right]$. Despite student logits $\left[0.4,0.4\right]$ and $\left[0.6,0.6\right]$ yielding equal losses under the mean square error metric, they convey different meanings. Specifically, the logit $\left[0.4,0.4\right]$ lies below the decision plane, while $\left[0.6,0.6\right]$ lies above it. This highlights potential overfitting in intra-sample distillation. To address this, additional constraints are necessary.
III-B2 Inter-Sample Distillation
Knowledge distillation should guarantee numerical similarity and semantic consistency between teacher logits and corresponding student logits [70, 4]. However, intra-sample alignment only emphasizes numerical similarity to align target logits. We introduce another inter-sample semantic alignment to simultaneously separate the logits of different samples. The inter-sample constraint $\mathcal{L}_{\text{inter}}$ is defined as:
$\mathcal{L}_{\text{inter}}$ $=-\frac{1}{n(n-1)}\sum_{i=0}^{n}\sum_{j\neq i}^{n}\left[d\left(\boldsymbol{s}_{i},\boldsymbol{s}_{j}\right)+d\left(\boldsymbol{s}_{i},\boldsymbol{t}_{j}\right)\right]$ (5)
$=-\underset{\begin{subarray}{c}{\boldsymbol{x}_{i}\sim\mathcal{X}},{\boldsymbol{x}_{j}\sim\mathcal{X}}\end{subarray}}{\mathbb{E}}{d\left(\boldsymbol{s}_{i},\boldsymbol{s}_{j}\right)}-\underset{\begin{subarray}{c}{\boldsymbol{x}_{i}\sim\mathcal{X}},{\boldsymbol{x}_{j}\sim\mathcal{X}}\end{subarray}}{\mathbb{E}}{d\left(\boldsymbol{s}_{i},\boldsymbol{t}_{j}\right)}.$
where $\boldsymbol{x}_{i}\neq\boldsymbol{x}_{j}$. The objective $\mathcal{L}_{\text{inter}}$ is to make $\boldsymbol{s}_{i}$ far from both $\boldsymbol{s}_{j}$ and $\boldsymbol{t}_{j}$. As shown in Fig. 3(b), we can alleviate the misalignment of $\boldsymbol{s}_{i}$ and $\boldsymbol{t}_{j}$ with Eq. 5 by maximizing their distance simultaneously.
Generally, inter-sample pairs include $\left(\boldsymbol{s}_{i},\boldsymbol{s}_{j}\right)$ and $\left(\boldsymbol{s}_{i},\boldsymbol{t}_{j}\right)$. We consider intra-sample pairs and can obtain triples $\left(\boldsymbol{s}_{i},\boldsymbol{t}_{i},\boldsymbol{s}_{j}\right)$ and $\left(\boldsymbol{s}_{i},\boldsymbol{t}_{i},\boldsymbol{t}_{j}\right)$. Fig. 4 visualizes these two triple categories. Previous knowledge distillation approaches [70, 4] use both triple types for training. However, unlike these methods [64, 21, 8] in which each triple element is trainable, the teacher model remains fixed, with its output $\boldsymbol{t}_{i}$ unchanged. It may be questionable to simply apply these two triples in knowledge distillation methods.
To address this issue, we only use triples $\left(\boldsymbol{s}_{i},\boldsymbol{t}_{i},\boldsymbol{s}_{j}\right)$ during training. Fig. 4 briefly illustrates the rationale behind this selection. In Fig. 4(a), $\boldsymbol{s}_{i}$ approaches $\boldsymbol{t}_{i}$ for intra-sample alignment while distancing itself from $\boldsymbol{s}_{j}$ for inter-sample alignment. Even with a gradient conflict when $\theta_{i}>\pi/2$, we can resolve it by adjusting $\boldsymbol{s}_{j}$. However, in Fig. 4(b), the conflict remains unresolved due to the fixed $\boldsymbol{t}_{j}$. Additionally, Fig. 4(c) depicts the worst-case scenario where the gradient conflict is at its maximum. From Fig. 4, we can conclude that inter-samples are crucial for model updates and the triple $\left(\boldsymbol{s}_{i},\boldsymbol{t}_{i},\boldsymbol{t}_{j}\right)$ may cause gradient conflicts. Thus, our method merely adopt the triple $\left(\boldsymbol{s}_{i},\boldsymbol{t}_{i},\boldsymbol{s}_{j}\right)$, and the inter-sample objective is formed as:
$\mathcal{L}_{\text{inter}}$ $=-\frac{1}{n(n-1)}\sum_{i=0}^{n}\sum_{j\neq i}^{n}d\left(\boldsymbol{s}_{i},\boldsymbol{s}_{j}\right)$ (6)
$=-\underset{\begin{subarray}{c}{\boldsymbol{x}_{i}\sim\mathcal{X}},{\boldsymbol{x}_{j}\sim\mathcal{X}}\end{subarray}}{\mathbb{E}}{d\left(\boldsymbol{s}_{i},\boldsymbol{s}_{j}\right)}.$
Finally, we combine Eq. 2 and Eq. 6 with a tradeoff parameter $\beta$ to obtain the following loss function:
$\mathcal{L}_{\text{KD}}=\mathcal{L}_{\text{intra}}+\beta\mathcal{L}_{\text{inter}}.$ (7)
Eq. 7 is to maximize similarities between teacher and student logits for the same sample while minimizing correlations across different samples. By pulling positive pairs closer and pushing negative pairs apart, the intra-sample and inter-sample formulations enable the alignment of teacher and student logits, ensuring they share similar semantic information for similarity and dissimilarity.
Figure 5: Pseudo code of CKD in a Numpy-like style.
III-B3 Contrastive Formulation
The proposed structure-preserving alignment shares a similar objective with contrastive learning [54, 79]. We seek to transform our objective into a sample-wise contrastive learning formulation. Specifically, $\mathcal{L}_{\text{KD}}$ can be expressed as:
$\begin{split}\mathcal{L}_{\text{KD}}&=\underset{\boldsymbol{x}_{i}\sim\mathcal{X}}{\mathbb{E}}{d\left(\boldsymbol{t}_{i},\boldsymbol{s}_{i}\right)}-\beta\underset{\begin{subarray}{c}{\boldsymbol{x}_{i}\sim\mathcal{X}}\\ {\boldsymbol{x}_{j}\sim\mathcal{X}}\end{subarray}}{\mathbb{E}}{d\left(\boldsymbol{s}_{i},\boldsymbol{s}_{j}\right)}.\end{split}$ (8)
For simplicity, we utilize the distance metric $f(\boldsymbol{u},\boldsymbol{v})$ to assess the similarity between $\boldsymbol{u}$ and $\boldsymbol{v}$, and set $f(\boldsymbol{u},\boldsymbol{v})=-d(\boldsymbol{u},\boldsymbol{v})$. The loss function can be derived as:
$\mathcal{L}_{\text{KD}}$ $=\underset{\boldsymbol{x}_{i}\sim\mathcal{X}}{\mathbb{E}}\left[-(f\left(\boldsymbol{t}_{i},\boldsymbol{s}_{i}\right)-\beta\underset{\boldsymbol{x}_{j}\sim\mathcal{X}}{\mathbb{E}}f\left(\boldsymbol{s}_{i},\boldsymbol{s}_{j}\right))\right]$ (9)
$=\underset{\boldsymbol{x}_{i}\sim\mathcal{X}}{\mathbb{E}}\ln\left[{\frac{\underset{\boldsymbol{x}_{j}\sim\mathcal{X}}{\mathbb{E}}e^{\beta f\left(\boldsymbol{s}_{i},\boldsymbol{s}_{j}\right)}}{e^{f\left(\boldsymbol{t}_{i},\boldsymbol{s}_{i}\right)}}}+\frac{e^{f\left(\boldsymbol{t}_{i},\boldsymbol{s}_{i}\right)}}{e^{f\left(\boldsymbol{t}_{i},\boldsymbol{s}_{i}\right)}}-1\right]$ (10)
$\simeq\underset{\boldsymbol{x}_{i}\sim\mathcal{X}}{\mathbb{E}}\left[-\ln{\frac{e^{f\left(\boldsymbol{t}_{i},\boldsymbol{s}_{i}\right)}}{e^{f\left(\boldsymbol{t}_{i},\boldsymbol{s}_{i}\right)}+\underset{\boldsymbol{x}_{j}\sim\mathcal{X}}{\mathbb{E}}e^{\beta f\left(\boldsymbol{s}_{i},\boldsymbol{s}_{j}\right)}}}\right]$ (11)
$\simeq\underset{\boldsymbol{x}_{i}\sim\mathcal{X}}{\mathbb{E}}\left[-\ln{\frac{e^{f\left(\boldsymbol{t}_{i},\boldsymbol{s}_{i}\right)/\tau}}{e^{f\left(\boldsymbol{t}_{i},\boldsymbol{s}_{i}\right)/\tau}+\underset{\boldsymbol{x}_{j}\sim\mathcal{X}}{\mathbb{E}}e^{\beta f\left(\boldsymbol{s}_{i},\boldsymbol{s}_{j}\right)/\tau}}}\right].$ (12)
The objective can be viewed as a sample-wise contrastive loss. Here, the teacher logit $\boldsymbol{t}_{i}$ and its corresponding student logit $\boldsymbol{s}_{i}$ form a positive pair, while $\boldsymbol{s}_{i}$ and $\boldsymbol{s}_{j}$ form a negative pair. Next, we will explore the benefits of this contrastive formulation.
(a) Traditional class-wise contrastive
(b) Typical class-wise contrastive: CRD
(c) Proposed sample-wise contrastive
Figure 6: Visualization of two categories of contrastive learning: Class-wise vs. Sample-wise. CRD is a typical class-wise formulation. Different background colors represent various sample categories, with P denoting positive samples and N denoting negative samples. The symbol E indicates that samples from the same class are excluded during training. The dashed circle contains teacher and student representations of the same sample
As shown in Eq. 9, to calculate the expected value $f(\boldsymbol{s}_{i},\boldsymbol{s}_{j})$ for each $\boldsymbol{x}_{i}$, we sample negative pairs from the data space $\boldsymbol{x}_{j}$. Following vanilla KD [25], we introduce a temperature parameter $\tau$ to soften the probability distribution, satisfying Eq. 12. The parameter $\beta$ balances the importance of positive and negative pairs, where a larger $\beta$ is assigned when negative pairs are important; otherwise, a smaller value is preferable. Note that negative sampling can also achieve this trade-off. When the negative space is significant, we increase the sampling of negative pairs for more detailed space estimation and weight adjustment. Conversely, fewer negative samples are selected when they are less important. Contrastive learning enables efficient batch data processing in parallel. In our formulation, we construct batches of training data, each containing one positive sample and $B-1$ negative samples, where $B$ is the batch size. From Fig. 2, the number of negative samples depends on $B$. Therefore, by varying $B$, we can sample different numbers of negative samples, balancing intra-sample and inter-sample alignment.
Additionally, we substitute $f\left(\boldsymbol{t}_{i},\boldsymbol{s}_{j}\right)$ for $f\left(\boldsymbol{s}_{i},\boldsymbol{s}_{j}\right)$. First, our method primarily aims to align $\boldsymbol{s}_{i}$ and $\boldsymbol{t}_{i}$, such that $\boldsymbol{t}_{i}=\boldsymbol{s}_{i}$. This leads to $\boldsymbol{t}_{i}=\boldsymbol{s}_{i}\Rightarrow f\left(\boldsymbol{t}_{i},\boldsymbol{s}_{j}\right)=f\left(\boldsymbol{s}_{i},\boldsymbol{s}_{j}\right)$. Then, $f\left(\boldsymbol{s}_{i},\boldsymbol{s}_{j}\right)$ involves two updating variables but $f\left(\boldsymbol{t}_{i},\boldsymbol{s}_{j}\right)$ only has one variable, simplifying the optimization of $\boldsymbol{s}_{j}$. Specifically, we treat $\boldsymbol{t}_{i}$ as an anchor logit and separately learn its positive logit $\boldsymbol{s}_{i}$ and negative logit $\boldsymbol{s}_{j}$. Most importantly, this setup allows our method to leverage negative samples through the N-pair loss [64] or InfoNCE loss [54]. Similar to CLIP [58], we sample a batch of $B$ (teacher, student) pairs with $B\times B$ (teacher, student) pairings for efficient training.
Consequently, we derive the final loss function as,
$\mathcal{L}_{\text{KD}}$ $\simeq\underset{\boldsymbol{x}_{i}\sim\mathcal{X}}{\mathbb{E}}\left[-\ln{\frac{e^{f\left(\boldsymbol{t}_{i},\boldsymbol{s}_{i}\right)/\tau}}{e^{f\left(\boldsymbol{t}_{i},\boldsymbol{s}_{i}\right)/\tau}+{\textstyle\sum\limits^{B}}e^{f\left(\boldsymbol{s}_{i},\boldsymbol{s}_{j}\right)/\tau}}}\right]$ (13)
$\simeq\underset{\boldsymbol{x}_{i}\sim\mathcal{X}}{\mathbb{E}}\left[-\ln{\frac{e^{f\left(\boldsymbol{t}_{i},\boldsymbol{s}_{i}\right)/\tau}}{e^{f\left(\boldsymbol{t}_{i},\boldsymbol{s}_{i}\right)/\tau}+{\textstyle\sum\limits^{B}}e^{f\left(\boldsymbol{t}_{i},\boldsymbol{s}_{j}\right)/\tau}}}\right].$ (14)
In Fig. 5, we present the pseudo-code for the proposed CKD in a Numpy-like style.
Theorem 1.
Given triples $\left(\boldsymbol{t}_{i},\boldsymbol{s}_{i},\boldsymbol{s}_{j}\right)$ and the loss function $\mathcal{L}_{i}$ that is defined as:
$\mathcal{L}_{i}=-\ln{\frac{e^{f\left(\boldsymbol{t}_{i},\boldsymbol{s}_{i}\right)/\tau}}{e^{f\left(\boldsymbol{t}_{i},\boldsymbol{s}_{i}\right)/\tau}+{\textstyle\sum\limits^{B}}e^{f\left(\boldsymbol{t}_{i},\boldsymbol{s}_{j}\right)/\tau}}}.$ (15)
The gradients $\nabla_{\boldsymbol{s}_{i}}\mathcal{L}_{i}$ and $\nabla_{\boldsymbol{s}_{j}}\mathcal{L}_{i}$ are proportional to $g_{i}$ that is formed as:
$g_{i}=1-\frac{e^{f\left(\boldsymbol{t}_{i},\boldsymbol{s}_{i}\right)/\tau}}{e^{f\left(\boldsymbol{t}_{i},\boldsymbol{s}_{i}\right)/\tau}+{\textstyle\sum\limits^{B}}e^{f\left(\boldsymbol{t}_{i},\boldsymbol{s}_{j}\right)/\tau}}.$ (16)
Proof.
The detailed proof can be found in [84]. ∎
Theorem 2.
For class-wise contrastive learning, $\boldsymbol{x}_{i}$ and $\boldsymbol{x}_{j}$ belong to different classes and $f\left(\boldsymbol{t}_{i},\boldsymbol{s}_{i}\right)-f\left(\boldsymbol{t}_{i},\boldsymbol{s}_{j}\right)=\rho_{j}$. When we decrease $B$, The gradients $\nabla_{\boldsymbol{s}_{i}}\mathcal{L}_{i}$ also decreases.
Proof.
From Eq. 16, we can obtain following equations:
$g_{i}$ $=1-\frac{e^{f\left(\boldsymbol{t}_{i},\boldsymbol{s}_{i}\right)/\tau}}{e^{f\left(\boldsymbol{t}_{i},\boldsymbol{s}_{i}\right)/\tau}+{\textstyle\sum\limits^{B}}e^{f\left(\boldsymbol{t}_{i},\boldsymbol{s}_{j}\right)/\tau}}$ (17)
$=1-\frac{1}{1+{\textstyle\sum\limits^{B}}e^{-\rho_{j}/\tau}}$ (18)
$\simeq 1-\frac{1}{1+B\times e^{-\rho_{j}/\tau}}.$ (19)
When $B$ decreases, we sample fewer negative samples, resulting in a larger $\rho_{j}$ and a smaller $g_{i}$. Consequently, the gradients $\nabla_{\boldsymbol{s}_{i}}\mathcal{L}_{i}$ decrease. ∎
From Theorem 2, we can find that when positive samples and negative samples exhibit weak correlations, the value of $f(\boldsymbol{t}_{i},\boldsymbol{s}_{j})$ is minimized, reducing the gradient of backpropagation $\nabla_{\boldsymbol{s}_{i}}\mathcal{L}_{i}$. This hampers model parameter updates. Therefore, class-wise methods [64, 21] heavily rely on large batch sizes to mitigate this issue and improve learning efficiency [84].
Input: Images $\mathcal{X}={\{\boldsymbol{x}_{i}\}_{i=1}^{n-1}}$, annotations $\mathcal{Y}={\{y_{i}\}_{i=1}^{n-1}}$, and the teacher model $\mathcal{T}(\mathcal{X};\theta_{\mathcal{T}})$
Output: The student model $\mathcal{S}(\mathcal{X};\theta_{\mathcal{S}})$
for epoch $\in\{1,2,\ldots,\text{Epoch}\}$ do
       while Randomly sample a batch $\boldsymbol{B}\in\mathcal{X}$ and $\boldsymbol{y}\in\mathcal{Y}$ do
             % Compute teacher and student logits
             $\boldsymbol{t}_{B}=\mathcal{T}(\boldsymbol{B};\theta_{\mathcal{T}})$
             $\boldsymbol{s}_{B}=\mathcal{S}(\boldsymbol{B};\theta_{\mathcal{S}})$
            % Compute knowledge distillation loss
             $\boldsymbol{t^{\prime}}_{B}=$ Normalize($\boldsymbol{t}_{B}$)
             $\boldsymbol{s^{\prime}}_{B}=$ Normalize($\boldsymbol{s}_{B}$)
             $\mathcal{L}_{\text{KD}}\leftarrow$ SampleWiseContrastive($\boldsymbol{s^{\prime}}_{B},\boldsymbol{t^{\prime}}_{B}$)
            % Compute classification loss
             $\boldsymbol{\bar{y}}\leftarrow$ Softmax($\boldsymbol{s}_{B}$)
             $\mathcal{L}_{\text{Task}}\leftarrow$ CrossEntropy($\boldsymbol{\bar{y}}$, $\boldsymbol{y}$)
            % Apply the loss in Eq. 7
             $\mathcal{L}\leftarrow\mathcal{L}_{\text{Task}}+\alpha\mathcal{L}_{\text{KD}}$
             $\theta_{\mathcal{S}}\leftarrow\theta_{\mathcal{S}}-\eta\ \partial\mathcal{L}/\partial\theta_{\mathcal{S}}$
       end while
end for
Algorithm 1 Training procedure of the proposed CKD
TABLE I: Classification accuracy (%) on the CIFAR-100 validation set, where teacher and student networks share the same architecture type. $\mathrm{\Delta}$ denotes the improved performance compared to vanilla KD. All results are the average over five trials.
Methods Teacher ResNet56 ResNet110 ResNet32$\times$4 WRN-40-2 WRN-40-2 VGG13
72.34 74.31 79.42 75.61 75.61 74.64
Student ResNet20 ResNet32 ResNet8$\times$4 WRN-16-2 WRN-40-1 VGG8
69.06 71.14 72.50 73.26 71.98 70.36
Features FitNet[59] 69.21 71 06 73.50 73.58 72.24 71.02
AT[31] 70.55 72.31 73.44 74.08 72.77 71.43
RKD[55] 69.61 71.82 71.90 73.35 72.22 71.48
CRD[70] 71.16 73.48 75.51 75.48 74.14 73.94
OFD[24] 70.98 73.23 74.95 75.24 74.33 73.95
ReviewKD[7] 71.89 73.89 75.63 76.12 75.09 74.84
Logits Vanilla KD[25] 70.66 73.08 73.33 74.92 73.54 72.98
TAKD[51] 70.83 73.37 73.81 75.12 73.78 73.23
DKD[90] 71.97 74.11 76.32 76.24 74.81 74.68
CKD 72.12 74.48 76.76 76.28 75.14 74.98
$\mathrm{\Delta}$ +1.46 +1.40 +3.43 +1.36 +1.60 +2.00
Theorem 3.
For sample-wise contrastive learning, $\exists j$, $\boldsymbol{x}_{i}$ and $\boldsymbol{x}_{j}$ are in the same class, such that $\exists j,\;f\left(\boldsymbol{t}_{i},\boldsymbol{s}_{i}\right)\sim f\left(\boldsymbol{t}_{i},\boldsymbol{s}_{j}\right)$, large batch sizes are not crucial.
Proof.
For sample-wise triples, $\exists j,f\left(\boldsymbol{t}_{i},\boldsymbol{s}_{i}\right)\sim f\left(\boldsymbol{t}_{i},\boldsymbol{s}_{j}\right)$, we can obtain $\exists j,\rho_{j}\sim 0$ and a larger $B\times e^{-\rho_{j}}$ is achieved, when we sample negative samples that are from the same class. Our method is unnecessary for comprehensive negative space sampling with large batch sizes. ∎
Fig. 6 visualizes traditional class-wise and the proposed sample-wise formulations. Specifically, CRD [70] is a representative class-wise contrastive learning approach by using a memory bank. We can observe that class-wise formulations including CRD generate negative pairs from different classes. These negative samples often have low similarity due to varied semantic content, necessitating a larger batch size to increase $g_{i}$. In contrast, the sample-wise formulation selects negative samples from the same class, thereby exhibiting stronger relationships. Thus, our method does not require a large batch size. In fact, a large batch size can hinder our method by influencing intra-sample alignment with a large $\beta$. However, a small batch size may also lead to insufficient gradient magnitude.
Specifically, given the logit space $\boldsymbol{T}\in\mathbb{R}^{c\times n}$, the negative logit space for the $i$-th sample is denoted as $\boldsymbol{T}_{/i}\in\mathbb{R}^{c\times(n-1)}$. Assume that there are $c$ classes, each containing $m$ samples. We can derive:
$\text{rank}\left(\boldsymbol{T}_{/i}\right)\leq\min{\left(c,n-1\right)}=c,\;\textit{ s.t. }c\ll n.$ (20)
Eq. 20 suggests that in the proposed method, the number of negative pairs sampled should heavily depend on the class number $c$ rather than the sample number $n$. Taking category correlations into consideration, the actual number of negative pairs sampled should be even lower than $c$.
Furthermore, compared to class-wise contrastive learning approaches, our method can be efficiently trained with well-designed triples. For classic contrastive learning methods with $n$ samples, a total of $\mathcal{O}(nm)$ triples should be selected for each sample:
$\binom{m-1}{1}\binom{n-m}{1}=nm-m^{2}-n+m\propto\mathcal{O}(nm)$ (21)
Eq. 21 shows that for each sample, we randomly select one sample within the same class including $m-1$ selections, and one sample from a different class, yielding a total of $\mathcal{O}(nm)$ triples. As for CRD, we only need to select one sample from different classes because the positive pair is obtained by using the same sample:
$\binom{1}{1}\binom{n-m}{1}=n-m\propto\mathcal{O}(n).$ (22)
However, the negative space $\mathcal{O}(n)$ is vast, necessitating the use of a memory bank with a size of $K$. As a result, $K$ triples are available for training each sample. In contrast, the proposed method requires sampling only $\mathcal{O}(c)$ triples:
$\binom{1}{1}\binom{c}{1}=c\propto\mathcal{O}(1),$ (23)
where $c\ll n$ We randomly shuffle data samples and select non-overlapping $c$ samples in order. Therefore, our sample-wise formulation exhibits efficient trainability in comparison to the class-wise formulation in contrastive knowledge distillation.
To sum up, we encapsulate the training procedure in Algorithm 1.
TABLE II: Classification accuracy (%) on the CIFAR-100 validation set for different teacher-student network architecture. $\mathrm{\Delta}$ denotes the improvement in performance compared to vanilla KD. Results are averaged over five trials.
Methods Teacher ResNet32$\times$4 WRN-40-2 VGG13 ResNet50 ResNet32$\times$4
79.42 75.61 74.64 79.34 79.42
Student ShuffleNetV1 ShuffleNetV1 MobileNetV2 MobileNetV2 ShuffleNetV2
70.50 70.50 64.60 64.60 71.82
Features FitNet[59] 73.59 73.73 64.14 63.16 73.54
AT[31] 71.73 73.32 59.40 58.58 72.73
RKD[55] 72.28 72.21 64.52 64.43 73.21
CRD[70] 75.11 76.05 69.73 69.11 75.65
OFD[24] 75.98 75.85 69.48 69.04 76.82
ReviewKD[7] 77.45 77.14 70.37 69.89 77.78
Logits Vanilla KD[25] 74.07 74.83 67.37 67.35 74.45
TAKD[51] 74.53 75.34 67.91 68.02 74.82
DKD[90] 76.45 76.70 69.71 70.35 77.07
CKD 77.72 77.47 70.11 70.62 78.18
$\mathrm{\Delta}$ +3.65 +2.64 +2.74 +3.27 +3.73
IV Experiments
IV-A Image Classification
IV-A1 Datasets
We first conduct experiments on CIFAR-100 [32] and ImageNet-1K [13]. The CIFAR-100 dataset contains 100 classes, each with 600 images. The training and validation sets comprise 50K and 10K images. The ImageNet-1K dataset consists of 1000 classes. The training set has about 1.28 million images, and the validation set contains 50k images.
TABLE III: Top-1 and Top-5 accuracy (%) on the ImageNet-1K validation set. ResNet-34 serves as the teacher, while ResNet-18 acts as the student. $\mathrm{\Delta}$ denotes the performance improvement compared to vanilla KD. All results are the average over three trials.
Baselines Features Logits
Teacher Student AT[85] OFD[24] CRD[70] ReviewKD[7] Vanilla KD[25] DKD[90] CKD $\mathrm{\Delta}$
Top-1 73.31 69.75 70.69 70.81 71.17 71.61 70.66 71.70 72.24 +1.58
Top-5 91.42 89.07 90.01 89.98 90.13 90.51 89.88 90.41 90.81 +0.93
TABLE IV: Top-1 and Top-5 accuracy (%) on the ImageNet-1K validation set. ResNet-34 serves as the teacher, while ResNet-18 acts as the student. $\mathrm{\Delta}$ denotes the improved performance compared to vanilla KD. All results are averaged over three trials.
Baselines Features Logits
Teacher Student AT[85] OFD[24] CRD[70] ReviewKD[7] Vanilla KD[25] DKD[90] CKD $\mathrm{\Delta}$
Top-1 76.16 68.87 69.56 71.25 71.37 72.56 68.58 72.05 72.97 +4.39
Top-5 92.86 88.76 89.33 90.34 90.41 91.00 88.98 91.05 91.36 +2.38
TABLE V: Object detection performance on the MS-COCO dataset. We employ the two-stage Faster RCNN with FPN as the detector. Teacher-student pairs are ResNet101 & ResNet-18, ResNet101 & ResNet50, and ResNet50 & MobileNetV2, respectively. $\mathrm{\Delta}$ represents the improved performance compared to ReviewKD.
Methods ResNet101 & ResNet18 ResNet101 & ResNet50 ResNet50 & MobileNetV2
AP AP50 AP75 AP AP50 AP75 AP AP50 AP75
Teacher 42.04 62.48 45.88 42.04 62.48 45.88 40.22 61.02 43.81
Student 33.26 53.61 35.26 37.93 58.84 41.05 29.47 48.87 30.90
Vanilla KD [25] 33.97 54.66 36.62 38.35 59.41 41.71 30.13 50.28 31.35
FitNet[59] 34.13 54.16 36.71 38.76 59.62 41.80 30.20 49.80 31.69
FGFI[76] 35.44 55.51 38.17 39.44 60.27 43.04 31.16 50.68 32.92
ReviewKD[7] 36.75 56.72 34.00 40.36 60.97 44.08 33.71 53.15 36.13
DKD[90] 35.05 56.60 37.54 39.25 60.90 42.73 32.34 53.77 34.01
DKD+ReviewKD 37.01 57.53 39.85 40.65 61.51 44.44 34.35 54.89 36.61
CKD 35.08 55.47 38.01 39.71 60.64 43.31 31.72 50.01 33.62
CKD+ReviewKD 37.65 57.91 40.71 41.45 61.93 45.31 35.31 55.35 37.88
$\mathrm{\Delta}$ +0.90 +1.19 +6.71 +1.09 +0.96 +1.23 +1.60 +2.20 +1.75
IV-A2 Experimental Setup
We evaluate eleven isomorphic or heterogeneous architectures on CIFAR-100 and compare with vanilla KD [25], AT [85], FitNet [59], RKD [55], CRD [70], OFD [24], ReviewKD [7], TAKD[51], and DKD [90].
For ImageNet-1K, we evaluate representatives, including ResNet34 & ResNet18, and ResNet50 & MobileNetV2. We provide comparisons with vanilla KD [25], AT [85], OFD [24], CRD [70], ReviewKD [7], and DKD [90]. Top-1 and top-5 recognition rates are reported.
Figure 7: Comparison of temperatures on CIFAR-100. Experiments are conducted on (1) ResNet8$\times$4 and ResNet32$\times$4, (2) ShuffleNetV1 and ResNet32$\times$4.
Figure 8: Comparison of batch sizes on CIFAR-100. Experiments are conducted on (1) ResNet8$\times$4 and ResNet32$\times$4, (2) ShuffleNetV1 and ResNet32$\times$4.
IV-A3 Implementation Details
All experiments are conducted on NVIDIA RTX 3090. For CIFAR-100, we set $\alpha$ to 100 and employ cosine annealing for learning rates. For homogeneous and heterogeneous networks, we set batch sizes to 64 and 32, with an initial learning rate of 0.05 and 0.01, respectively. For ImageNet-1K, we also use cosine annealing, with an initial learning rate of 0.2. We set $\alpha$ to 10 and a batch size to 512.
Figure 9: Comparison of training time per batch and accuracy on CIFAR-100. The teacher model is ResNet32$\times$4 and the student model is ResNet8$\times$4. Additional parameters are denoted after $+$, which are added based on the student’s foundation.
TABLE VI: Ablation studies of contrastive formulation on the CIFAR-100 dataset. Experiments are conducted on homogeneous and heterogeneous teacher-student pairs. SupCon and CRD are class-wise methods, while our CKD is a sample-wise approach.
Methods Homogeneous architectures Heterogeneous architectures
Teacher ResNet56 WRN-40-2 WRN-40-2 VGG13 ResNet50 ResNet32$\times$4
72.34 75.61 75.61 74.64 79.34 79.42
Student ResNet20 WRN-16-2 WRN-40-1 MobileNetV2 MobileNetV2 ShuffleNetV2
69.06 73.26 71.98 64.60 64.60 71.82
SupCon[28] 69.93 75.45 74.10 67.39 67.72 75.20
CRD[70] 71.16 75.48 74.14 69.73 69.11 75.65
CKD 72.12 76.28 75.14 70.11 70.62 78.18
TABLE VII: Ablation studies of triple design on the CIFAR-100 dataset. Experiments are conducted on both homogeneous and heterogeneous teacher-student pairs. Due to the fixed teacher, triples $\left(\mathbf{t}_{i},\mathbf{s}_{i},\mathbf{t}_{j}\right)$ is meaningless and excluded from this table.
Triples Homogeneous architectures Heterogeneous architectures
Teacher ResNet56 WRN-40-2 WRN-40-2 VGG13 ResNet50 ResNet32$\times$4
72.34 75.61 75.61 74.64 79.34 79.42
Student ResNet20 WRN-16-2 WRN-40-1 MobileNetV2 MobileNetV2 ShuffleNetV2
69.06 73.26 71.98 64.60 64.60 71.82
CKD$\left(\mathbf{s}_{i},\mathbf{t}_{i},\mathbf{s}_{j}\right)$ 71.76 75.31 74.40 70.04 68.93 78.02
CKD$\left(\mathbf{s}_{i},\mathbf{t}_{i},\mathbf{t}_{j}\right)$ 71.69 75.88 74.43 70.09 69.68 77.06
CKD$\left(\mathbf{t}_{i},\mathbf{s}_{i},\mathbf{s}_{j}\right)$ 72.12 76.28 75.14 70.11 70.62 78.18
IV-A4 Performance on CIFAR-100
We conduct classification experiments by using eleven combinations of isomorphic or heterogeneous teacher and student network architectures, including ResNet [22], WideResNet [73], VGG [63], ShuffleNet [88, 47], and MobileNet [60].
Table I presents classification performance by using teacher and student architectures of the same style. We can observe that our method achieves competitive performance against feature-based and logit-based methods. Specifically, the proposed CKD improves vanilla KD accuracy by 1.36%, with a maximum gain of 3.43%. Moreover, CKD also outperforms CRD on the CIFAR-100 dataset, with an improvement of almost 1.00%.
Table II shows the classification performance of teacher-student networks with different styles. We see that our CKD method consistently yields improvements across teacher-student network pairs compared to vanilla KD. Specifically, for different architectures, CKD achieves an average improvement of 3.21% over vanilla KD, with an impressive maximum gain of 3.73%.
IV-A5 Performance on ImageNet-1K
To further verify the ability of CKD, we perform experiments on the ImageNet-1K dataset. Table III and Table IV report experimental results on the ImageNet-1K dataset. We can see that our method achieves superior performance with both teacher-student architectures. Moreover, the performance of CKD surpasses state-of-the-art results, providing further evidence of the effectiveness of our approach. Specifically, the proposed CKD achieves accuracy improvements of 1.58% and 4.39% over vanilla KD with the same and different teacher-student architectures.
IV-B Object Detection
IV-B1 Datasets
We conduct experiments on the MS-COCO dataset [43]. MS-COCO is a popular dataset featuring 80 categories, divided into two subsets, including a training set of 118,000 images and a validation set of 5,000 images.
IV-B2 Experimental Setup
We analyze ResNet-101 & ResNet-18, ResNet-101 & ResNet-50, and ResNet-50 & MobileNetV2 teacher-student pairs. Moreover, we evaluate the MS-COCO validation set with the baseline provided by Detectron2.
IV-B3 Implementation Details
Experiments are performed on NVIDIA RTX 3090. We utilize Faster R-CNN with FPN as the feature extractor and implement CKD on the R-CNN head. We set $\alpha$ to 0.1 and the temperature $\tau$ to 1.0. Schedulers and task-specific loss settings are following Detectron2.
IV-B4 Performance on MS-COCO
Table V presents the object detection results on the MS-COCO dataset under three teacher-student settings. We can see that the proposed approach obtains comparable performance against logit-based approaches but inferior performance compared to feature-based approaches, particularly for ReviewKD. Object detection relies on spatial representations to locate objects. However, logits lack detailed spatial information [36, 76]. ReviewKD [7] conducts multi-level feature distillation, resulting in better performance. We combine ReviewKD with logit-based approaches to utilize spatial information. Notably, the combined approaches, including CKD+ReviewKD and DKD+ReviewKD, significantly improve object detection results compared to previous counterparts.
IV-C Ablation Study
IV-C1 Effects of Temperature $\tau$
In the training phase, a high $\tau$ yields a smooth distribution across classes, making distillation more challenging. But, a low $\tau$ leads to a sharp distribution. To investigate the effect of $\tau$ on CKD, we conduct ablation studies by using ResNet32×$\times 4$ ResNet8×$\times 4$ and ResNet32×$\times 4$ ShuffleNetV1 pairs. Typically, $\tau$ is set to a low value within a range of 0.07 to 4.0, while other parameters are kept constant.
TABLE VIII: Classification accuracy (%) on the CIFAR-100 validation set. WRN-16-2 is used as a student. In the left-hand section of the table, WRN series networks serve as teachers, while in the right-hand section, different series networks are utilized as teachers.
Methods WRN Series Different Series
Teacher WRN-40-2 WRN-16-4 WRN-28-4 VGG13 WRN-16-4 ResNet50
75.61 77.21 79.12 74.64 77.21 79.34
Studnet 73.26 73.26 73.26 73.26 73.26 73.26
vanilla KD[25] 74.92 76.00 75.14 74.30 76.00 74.62
CKD 76.28 76.22 76.42 75.14 76.22 76.45
Fig. 7 presents Top-1 classification accuracies for different temperatures $\tau$. We can observe that our method performs best when temperature $\tau$ is around 1.0 and performs poorly when $\tau$ is low or high. This suggests that our method is better suited for aligning ${t}_{i}$ and ${s}_{i}$ when the value of $\tau$ is near 1.0.
From Theorem 1, we conclude that the gradient $\nabla_{\boldsymbol{s}_{*}}\mathcal{L}_{i}$ is proportional to $g_{i}$. When $\tau<1$, the difference between $e^{f\left(\boldsymbol{t}_{i},\boldsymbol{s}_{i}\right)/\tau}$ and $e^{f\left(\boldsymbol{t}_{i},\boldsymbol{s}_{j}\right)/\tau}$ is magnified exponentially. In the limit, we obtain $\lim_{\tau\to 0}g_{i}=0$. As a result, $g_{i}$ approaches zero, causing a rapid decline in performance. When $\tau>1$, the difference is lessened and $g_{i}$ becomes more sensitive to negative samples as $\lim_{\tau\to\infty}g_{i}=\frac{B}{B+1}$.
IV-C2 Effects of Batchsize
To analyze the effect of batch size, we utilize two teacher-student pairs including ResNet56 & ResNet20 and ResNet32×$\times 4$ ShuffleNetV2. Moreover, we set the batch size to $2^{k}$, where $k\in[4,11]\cap k\in\mathbb{N}$.
Fig. 8 presents Top-1 classification performance with varying batch sizes. We can see that our method attains good performance when the batch size is 64 for ResNet56 & ResNet20 and 32 for ResNet32×$\times 4$ ShuffleNetV2. Larger or smaller batch sizes are not optimal.
We also find that a large batch size potentially hampers the efficacy of our method as it affects intra-sample alignment through an increased $\beta$. However, a small batch size can lead to insufficient gradient magnitude. In fact, the optimal batch size for the sample-wise formulation is strongly correlated with the number of categories. It should be smaller than the number of categories to avoid neglecting intra-sample distillation and facilitate effective inter-sample distillation.
Figure 10: Class Activation Map via GradCAM. The teacher networks (ResNet34) and the student networks (ResNet18) are trained on ImageNet-1K.
IV-C3 Influence of Contrastive Formulation
Fig. 6 visualizes class-wise and sample-wise contrastive formulations. SupCon [28] is a class-wise approach that selects positive pairs within the same class. Noat that CRD [70] is a typical class-wise method with positive pairs from the same sample and negative pairs from different classes. In contrast, CKD chooses negative pairs from different samples but potentially from the same class.
Table VI presents classification performance under different formulations. We can observe that CKD outperforms SupCon and CRD because it uses sample-wise negative samples. Moreover, CRD exploits a memory bank to access a large number of negative samples and achieves a higher accuracy than SupCon.
IV-C4 Influence of Triple Design
In knowledge distillation, the element of triples can be sampled from teacher or student models. Hence, there are several kinds of triples, including $\left(\boldsymbol{s}_{i},\boldsymbol{t}_{i},\boldsymbol{s}_{j}\right)$, $\left(\boldsymbol{s}_{i},\boldsymbol{t}_{i},\boldsymbol{t}_{j}\right)$, $\left(\boldsymbol{t}_{i},\boldsymbol{s}_{i},\boldsymbol{t}_{j}\right)$ and $\left(\boldsymbol{t}_{i},\boldsymbol{s}_{i},\boldsymbol{s}_{j}\right)$. $\left(\boldsymbol{s}_{i},\boldsymbol{t}_{i},\boldsymbol{s}_{j}\right)$ means that we align $\mathbf{s}_{i}$ and $\boldsymbol{t}_{i}$ while distancing $\boldsymbol{s}_{i}$ from $\boldsymbol{s}_{j}$. This triple can be implemented by modifying diagonal elements of the affinity matrix with $\boldsymbol{s}_{i}$ and $\boldsymbol{t}_{i}$. Due to fixed teacher logits, $\left(\boldsymbol{t}_{i},\boldsymbol{s}_{i},\boldsymbol{t}_{j}\right)$ equals intra-sample alignment.
Table VII shows the classification performance with different triples. We can see that the utilization of triples $\left(\boldsymbol{t}_{i},\boldsymbol{s}_{i},\boldsymbol{s}_{j}\right)$ yields better performance than the other triples. Our method anchors on $\boldsymbol{t}_{i}$, pulling $\boldsymbol{s}_{i}$ closer while pushing $\boldsymbol{s}_{j}$ apart. This objective is clear and simple. However, as discussed in Section III-B, triples $\left(\boldsymbol{s}_{i},\boldsymbol{t}_{i},\boldsymbol{t}_{j}\right)$ may cause gradient conflict. For triples $\left(\boldsymbol{s}_{i},\boldsymbol{t}_{i},\boldsymbol{s}_{j}\right)$, this strategy mainly performs contrastive learning in the student logits space.
IV-D More Analyses
IV-D1 Training Efficiency
To evaluate the training efficiency of our approach, we compare its accuracy and parameter count against state-of-the-art methods on the CIFAR-100 dataset. Fig. 9 depicts training time per batch and classification performance. We can observe that our method exhibits high training efficiency, striking a balance between model parameters and classification performance. Typically, feature-based methods require more parameters and longer training times than logit-based methods. ReviewKD and CRD, in particular, take twice the training time compared to our method, due to multi-level feature review and memory bank usage.
IV-D2 Effects of Teacher Capacity
Previous works [10, 74] have found that large teacher models may not be optimal for smaller student models due to disparities in parameters and capacity. Therefore, we further conduct experiments to analyze the influence of teacher networks with different capacities.
Table VIII presents classification performance on the CIFAR-100 dataset. We can see that our method achieves comparable performance across different WRN models, whereas vanilla KD exhibits variation in performance. Wide Residual Networks [73], is a ResNet variant. Therefore, our method also obtains an accuracy on par with ResNet50 as the teacher model, while vanilla KD has the second-worst performance. For vanilla KD, larger models lead to performance degeneration, for example, the accuracies of WRN-16-4 and WRN-28-4 are 76.00 and 75.14, respectively. Both the proposed method and vanilla KD perform poorly with the VGG13 architecture. Ablation studies prove that our method can partially mitigate the parameter and capacity gaps of larger models.
IV-E Visualizations
GradCAM identifies image regions critical for concept prediction. To evaluate the consistency of CKD with teacher networks, we employ GradCAM to visualize class activation maps in Fig. 10. The teacher-student pair, including ResNet34 and ResNet18, are trained on the ImageNet-1K dataset. We can observe that our method effectively handles complex scenes and multiple objects. For example, in the seventh column, two birds are present. Our method accurately replicates teacher knowledge by focusing on both birds, whereas student networks and vanilla KD focus only on the bird closer to the camera.
V Conclusion
In this paper, we have presented a contrastive knowledge distillation approach for image classification and object detection. To be specific, we formulate knowledge distillation as a logit alignment problem, mimicking the intra-sample numerical value and inter-sample semantic structure of teacher models. By converting the structure-preserving alignment into a sample-wise contrastive learning framework, our approach can be effectively and efficiently trained with well-designed positive and negative pairs. Experimental results demonstrate that our method achieves highly competitive performance against state-of-the-art methods on benchmark datasets. In the future, we will attempt to apply the proposed method to distill large models.

### v2

@@
 Abstract
-In this paper, we present a simple yet effective contrastive knowledge distillation approach, which can be formulated as a sample-wise alignment problem with intra- and inter-sample constraints.
-Unlike traditional knowledge distillation methods that concentrate on maximizing feature similarities or preserving class-wise semantic correlations between teacher and student features, our method attempts to recover the “dark knowledge” by aligning sample-wise teacher and student logits.
-Specifically, our method first minimizes logit differences within the same sample by considering their numerical values, thus preserving intra-sample similarities.
-Next, we bridge semantic disparities by leveraging dissimilarities across different samples.
-Note that constraints on intra-sample similarities and inter-sample dissimilarities can be efficiently and effectively reformulated into a contrastive learning framework with newly designed positive and negative pairs.
-The positive pair consists of the teacher’s and student’s logits derived from an identical sample, while the negative pairs are formed by using logits from different samples.
-With this formulation, our method benefits from the simplicity and efficiency of contrastive learning through the optimization of InfoNCE, yielding a run-time complexity that is far less than $\mathcal{O}(n^{2})$, where $n$ represents the total number of training samples.
-Furthermore, our method can eliminate the need for hyperparameter tuning, particularly related to temperature parameters and large batch sizes.
-We conduct comprehensive experiments on three datasets including CIFAR-100, ImageNet-1K, and MS COCO.
-Experimental results clearly confirm the effectiveness of the proposed method on both image classification and object detection tasks.
-Our source codes will be publicly available at https://github.com/wencheng-zhu/CKD.
+In this paper, we propose a simple yet effective contrastive knowledge distillation framework that achieves sample-wise logit alignment while preserving semantic consistency.
+Conventional knowledge distillation approaches exhibit over-reliance on feature similarity per sample, which risks overfitting, and contrastive approaches focus on inter-class discrimination at the expense of intra-sample semantic relationships.
+Our approach transfers ”dark knowledge” through teacher-student contrastive alignment at the sample level.
+Specifically, our method first enforces intra-sample alignment by directly minimizing teacher-student logit discrepancies within individual samples.
+Then, we utilize inter-sample contrasts to preserve semantic dissimilarities across samples.
+By redefining positive pairs as aligned teacher-student logits from identical samples and negative pairs as cross-sample logit combinations, we reformulate these dual constraints into an InfoNCE loss framework, reducing computational complexity lower than sample squares while eliminating dependencies on temperature parameters and large batch sizes.
+We conduct comprehensive experiments across three benchmark datasets, including the CIFAR-100, ImageNet-1K, and MS COCO datasets, and experimental results clearly confirm the effectiveness of the proposed method on image classification, object detection, and instance segmentation tasks.
 Index Terms:
-Knowledge distillation, contrastive learning, intra- and inter-sample constraint, sample-wise alignment
+Knowledge distillation, contrastive learning, intra-sample consistency, inter-sample contrast, logit alignment
 I Introduction
-Knowledge distillation (KD) [68, 15, 14] compresses knowledge without significant performance degradation, making it possible to deploy resource-intensive models on lightweight devices [40, 74, 19].
-Given the practical benefits, knowledge distillation has garnered much attention over the past decade, and numerous knowledge distillation approaches have been proposed and further introduced into a wide spectrum of tasks [10, 5, 75], such as image classification [39, 16, 71], object detection [87, 82, 61], and segmentation [45, 62, 81].
-While considerable improvements have been achieved, existing knowledge distillation methods still struggle to determine what knowledge to learn and how to learn it [66, 2, 53].
-Here, we posit that the student model can effectively replicate the output of the teacher model when each sample is appropriately aligned.
-Figure 1: Illustration of three typical approaches, including classic KD [25], CRD [70], and our CKD.
-Classic KD emphasizes feature similarities between teacher and student features, while CRD focuses on class-wise semantic information.
-In contrast, the proposed CKD conducts similarity alignment and captures sample-wise semantic structures simultaneously.
+Knowledge distillation (KD) [70, 15] effectively compresses knowledge without significant performance degradation, making it possible to deploy complex models on resource-limited devices [40, 76, 19].
+Owing to its practical benefits, knowledge distillation has garnered much attention over the past decade [10, 39].
+Numerous knowledge distillation approaches have been proposed [14, 89] and introduced into a wide spectrum of tasks [5, 77], such as image classification [16, 73], object detection [84, 63], and segmentation [45, 64, 83].
+Despite considerable improvements achieved, existing knowledge distillation methods still struggle to determine what knowledge to distill and how to distill it [68, 2, 54].
+Here, we posit that the student model can well replicate the outputs of the teacher model when each sample is properly aligned.
+Figure 1: Comparison of three knowledge distillation approaches.
+Classic KD aligns feature similarities between paired teacher-student samples, while CRD leverages contrastive learning to align class-wise semantics.
+In contrast, the proposed CKD preserves sample-wise similarity for intra-sample alignment while capturing structural semantics for inter-sample contrast.
 Hinton et al.
-[25] propose a seminal work on knowledge distillation, with the goal of reducing the Kullback-Leibler (KL) divergence between teacher and student logits.
-This influential research has led to a surge in related studies, most of which can be categorized into logit-based and feature-based approaches.
-Logit-based approaches distill the output logits of teachers [83, 35].
-Representative methods in this category are vanilla KD [25], DML [89], DIST [26], DKD [90], etc.
-Following vanilla KD [25], there has been a lack of research on logit distillation, mainly because of its inferior performance relative to feature distillation.
-However, logits contain compact information that is easier to distill.
-Furthermore, logit-based methods can take advantage of low computational and storage requirements [34, 37, 86].
-Our approach also employs logit distillation.
-Feature-based approaches seek to replicate features extracted from intermediate layers of a teacher model [42, 20].
-Typical methods include FitNet [59], OFD [24], ReviewKD [7], etc.
-These methods have advanced the field of knowledge distillation.
-However, due to rote memorization, they often suffer from severe overfitting that leads to low training losses but high test losses.
-Given that, RKD [55] resembles high-order structural relations among features instead of feature similarities.
-SP [72] preserves pairwise similarities in representation space.
-Although feature-based approaches effectively capture feature similarity structure, they may somewhat overlook semantic knowledge.
-CRD [70] innovatively distills structural representations through contrastive learning.
-To handle the requirement for large batch sizes in contrastive learning, CRD implements a memory buffer to enable the sampling of many negative samples, resulting in less-than-ideal training efficiency, particularly for dense prediction tasks.
-To address the aforementioned structure-preserving issues, we present a Contrastive Knowledge Distillation approach, dubbed CKD, which treats knowledge distillation as an intra- and inter-sample alignment problem.
+[25] introduced seminal work on knowledge distillation that transfers knowledge by minimizing the Kullback-Leibler (KL) divergence between teacher and student logits.
+Subsequent research has evolved along two main groups, including logit-based and feature-based approaches.
+Logit-based approaches mimic output teacher logits [85, 35].
+Representative methods are vanilla KD [25], DML [91], DIST [26], and DKD [92].
+Despite minimal computational and storage requirements [34, 37, 88], these approaches generally underperform feature-based methods.
+Feature-based approaches align intermediate layer representations between teacher and student networks [42, 20].
+Typical methods include FitNet [61], OFD [24], and ReviewKD [7].
+These methods achieve significant performance, however, they often overfit due to rote memorization, achieving a low training loss but poor generalization.
+Subsequent methods like RKD [56] address this limitation by modeling high-order feature correlations, whereas SP [74] enhances robustness through pairwise similarity preservation in representation space.
+While these techniques alleviate over-fitting to some extent, their core focus on structural relationships may neglect the transfer of semantic knowledge.
+Contrastive learning methods, such as CRD [72], enhance structural representation distillation.
+However, this approach encounters practical limitations: the reliance on large negative sample batches necessitates a large memory buffer, incurring high computational costs, especially for dense prediction tasks.
+To address structure-preserving limitations, we propose a Contrastive Knowledge Distillation approach, dubbed CKD, which unifies intra- and inter-sample distillation in a contrastive learning framework.
 Fig.
-1 briefly illustrates the main idea of the proposed approach in comparison with two representative approaches.
-We clearly observe that the proposed approach pulls logits from the same sample together while pushing logits from different samples apart.
-Specifically, our method assumes that a student model can well mimic a teacher model by distilling sample-wise information.
-This idea is straightforward and reasonable.
-On the one hand, previous works [59, 7] have shown tremendous success by maximizing intra-sample feature similarities.
-Inspired by these methods, we directly minimize logit differences within the same sample.
-On the other hand, solely relying on an intra-sample similarity constraint can easily lead to overfitting [57, 77].
-To this end, we dive into the inter-sample constraints among different samples for additional information.
-Finally, we integrate these two constraints into the loss function.
-Formally, the intra- and inter-sample constraints can be cast as a sample-wise contrastive learning problem.
-Compared to existing contrastive learning approaches [70, 4] that preserve the class-wise semantic structure, the sample-wise contrastive framework can distill teacher knowledge in an unsupervised way without label information.
-Moreover, our method precisely aligns teacher and student logits by exploiting both similarities and dissimilarities within each sample rather than each class.
-From the sample-wise perspective, this leads to a relatively small number of associated negative samples, without resorting to large batch sizes for sampling the whole negative space.
-Thus, the sample-wise formulation enables our approach to maintain low computational complexity in the absence of class-wise information.
-Furthermore, our method employs logit distillation over feature distillation for two main reasons.
-One reason is that logits contain high-level and compact semantic information, which is particularly suitable for contrastive learning.
-The other reason is that the dimension of logits, crucial for negative sampling, is lower.
-Notably, our method is model-agnostic and can be applied to feature distillation methods.
-To sum up, the contributions of this paper are summarized into three aspects:
+1 shows the main idea of the proposed approach compared to two representative approaches.
+We can observe that our approach aligns same-sample logits while separating cross-sample logits.
+Our method draws inspiration from the cross-modal alignment paradigm of CLIP, where we take knowledge distillation as a dual-modality alignment task between teacher and student embedding space with the fixed teacher networks.
+Unlike conventional approaches [61, 7] that enforce feature mimicking, our framework conducts cross-modal alignment through contrastive training, and this paradigm enables the student to capture the teacher’s semantic representations beyond feature consistency while leveraging the efficient sample utilization strategy of CLIP for distillation effectiveness [58, 79].
+However, directly applying CLIP is problematic as the teacher model is fixed.
+Therefore, we propose sample-wise contrastive learning with dual constraints: the intra-sample distillation enhances logit similarities within individual samples, and the inter-sample distillation increases differences across samples.
+The intra- and inter-sample alignment constraints can be formulated as a sample-wise contrastive learning problem.
+Unlike existing contrastive distillation methods [72, 4] that focus on class-wise semantic structures, the sample-wise framework has two key advantages.
+First, our method directly optimizes the similarity and dissimilarity of logits within individual sample rather than aligning teacher and student models at the class level.
+Second, by eliminating the dependency on class annotations, the framework supports fully unsupervised distillation.
+Notably, a key benefit of this sample-wise framework lies in its computational efficiency.
+Different from class-wise methods requiring large batch sizes to sample the entire negative space, our approach needs a minimal number of negative samples.
+Additionally, we prioritize logit distillation over feature distillation for two reasons.
+On one hand, logits provide compact and high-level semantic representations that are suitable for contrastive learning.
+On the other hand, compared to features, the lower dimensionality of logits significantly simplifies negative sampling.
+The framework remains model-agnostic and can be seamlessly integrated with feature distillation approaches.
+Our contributions can be summarized into three aspects:
   1. 1.
-    We present a contrastive knowledge distillation approach from the sample-wise perspective.
-Our method simultaneously conducts intra-sample similarity and inter-sample dissimilarity alignment between teacher and student logits.
+    We propose a contrastive knowledge distillation framework with dual constraints, in which our method simultaneously conducts intra-sample and inter-sample distillation between teacher and student logits.
   2. 2.
-    We formulate these two intra-sample and inter-sample constraints as a contrastive learning problem with well-designed positive and negative pairs.
-By using the proposed contrastive formulation, our method can be efficiently and effectively trained.
+    We cast these dual intra-sample and inter-sample constraints into a sample-wise contrastive formulation through customized positive and negative pairs, achieving effective and computationally efficient training.
   3. 3.
-    We evaluate the proposed approach on benchmarks for image classification and object detection.
-Without bells and whistles, our approach improves the accuracy of vanilla KD by more than 1.4% and 0.98% both on the CIFAR-100 and ImageNet-1K datasets.
+    Experiments on benchmark datasets validate the superiority of our method.
+Without bells and whistles, our method outperforms vanilla KD by 1.4% and 0.98% on the CIFAR-100 and ImageNet-1K datasets, respectively.
 II Related Work
@@
 II-A Knowledge Distillation
-Knowledge distillation is a representative model compression technique [65, 49, 80] that enables small models to attain strong performance of large models [1, 52, 48].
+Knowledge distillation is a widely adopted model compression technique [67, 50, 82] that allows compact models to achieve competitive performance with their larger counterparts [1, 53, 49].
 As a pioneer, Hinton et al.
-[25] introduced the concept of knowledge distillation that transfers “dark knowledge” by minimizing the KL divergence between teacher logits and student logits.
+[25] introduced the concept of knowledge distillation that transfers “dark knowledge” through minimization of the KL divergence between teacher and student logit distributions.
 Over the past few years, a rich line of knowledge distillation methods has been proposed [30, 11, 12].
-These knowledge distillation methods come in two flavors: logit distillation and feature distillation [50, 33].
+These methods come in two flavors: logit distillation and feature distillation [51, 33].
 II-A1 Logit Distillation
-Earlier logit distillation approaches attempt to enhance generalization performance through regularization and optimization techniques.
-However, there has been limited research on early-stage logit distillation due to inferior performance.
-DML [89] improved the generalization abilities of networks by collaboratively training an ensemble of models through mutual learning.
-Instead, TAKD [51] employed multi-step distillation via intermediate-sized networks, effectively bridging the knowledge transfer gap between teacher and student networks.
-ICKD-C [44] utilized grid-level inter-channel correlation to capture the intrinsic distribution and diversity of feature representations.
-GLD [29] employed a local spatial pooling layer for the simultaneous extraction of finely localized knowledge and holistic representations.
-DIST [26] leveraged both inter-class and intra-class relations to align probabilistic distributions and reinforce semantic similarities, respectively.
-DKD [90] decoupled logits into a target class and all non-target classes, thereby enhancing both the effectiveness and flexibility of knowledge transfer.
-Auto-KD [35] firstly introduced an automated search strategy by employing Monte Carlo tree search to optimize knowledge distillation architectures.
-CTKD [39] modulated the task difficulty levels of a student network by using a dynamic and learnable temperature setting.
-MLKD [27] also proposed a multi-level logit distillation framework, which distills instance-level, batch-level, and class-level information.
-LSKD [67] considered temperature as the standard deviation of logit and applied logit standardization before the softmax prediction.
-Besides, there are several other studies [57, 9] devoted to explaining the basic principle of knowledge distillation.
-Our method belongs to logit distillation that leverages contrastive information between the logits of the teacher and student models.
-Figure 2: The overall architecture of the proposed CKD.
-The newly designed triplets are introduced for similarity and semantic alignment.
-For example, we classify all other samples as negative, regardless of their class.
-In this formulation, training efficiency can be improved by reusing negative samples.
+Early logit distillation approaches mainly focused on enhancing model generalization through regularization and optimization.
+However, research on early-stage logit distillation has remained limited due to its inferior performance.
+Initial attempts like DML [91] improved network generalization via mutual learning across model ensembles, while TAKD [52] bridged the teacher-student capacity gap through multi-step distillation with intermediate networks.
+Subsequent advances introduced feature representation enhancement, distributional alignment, and automated framework design.
+ICKD-C [44] captured feature diversity through grid-level inter-channel correlations.
+GLD [29]combined local spatial pooling for fine-grained knowledge extraction.
+DIST [26] aligned inter-class and intra-class probabilistic relations, respectively.
+DKD [92] decoupled logits into a target class and all non-target classes to improve transfer flexibility.
+Auto-KD [35] employed a Monte Carlo tree search for architecture optimization.
+CTKD [39] dynamically adjusted task difficulty via learnable temperature scheduling.
+MLKD [27] proposed a multi-level logit distillation framework that distills instance-level, batch-level, and class-level information.
+LSKD [69] took temperature as the standard deviation of logit and applied logit standardization before the softmax prediction.
+Moreover, there are theoretical studies [58, 9] that further elucidate the principles of distillation.
+Our method belongs to logit distillation that leverages contrastive teacher-student logits to enhance knowledge transfer.
+Figure 2: The overall architecture of the proposed Contrastive Knowledge Distillation framework.
+The framework incorporates newly designed triplets to optimize intra-sample feature similarity and cross-sample semantics simultaneously.
+For each input instance, the teacher and student logits form positive pairs, while other samples within the mini-batch serve as negative pairs.
+This formulation significantly enhances training efficiency through computationally efficient negative sample reuse, effectively addressing the memory constraints typically associated with contrastive learning approaches.
 II-A2 Feature Distillation
-Feature distillation emphasizes the maximization of similarities between intermediate representations.
-FitNet [59] proposed a feature-based approach that extracted trainable student representations to predict teacher representations.
-But, the excessive proximity between samples often results in overfitting.
-Hence, AT [31] mimicked attention maps of a teacher model to enhance student performance.
-Instead, RKD [55] transferred the relational structure among data examples by incorporating both distance-wise and angle-wise distillation objectives.
-IRG [46] distilled not just instance features but also the often-overlooked instance relationships and cross-layer feature transformations.
-CCKD [56] captured both instance-level information and correlations between multiple instances for knowledge transfer.
-SP [72] introduced a similarity-preserving loss to maintain the activation pattern similarities of input pairs.
-OFD [24] combined a new margin ReLU, an optimized distillation feature position, and a selective L2 distance to enhance network compression.
-CRD [70] was the first contrastive distillation approach that preserves structural representations among features.
-ReviewKD [7] distilled feature representations from cross-level connection paths to guide the learning of a student model.
+Feature distillation focuses on aligning intermediate representations of teacher and student models.
+FitNet [61] pioneered feature distillation by using hint layers to match student features to teacher activations, but its strict alignment often caused overfitting.
+Subsequent methods adopted different alignment strategies.
+For example, AT [31] mimicked spatial attention maps of a teacher model to emphasize task-relevant regions.
+Differently, RKD [56] preserved relational structures between samples by using distance-wise and angle-wise distillation objectives.
+IRG [46] incorporated both instance relationships and cross-layer feature transformations.
+CCKD [57] captured multi-instance correlations for knowledge transfer.
+SP [74] introduced a similarity-preserving loss to maintain the activation patterns of input pairs.
+Instead, OFD [24] introduced a margin ReLU, an optimized distillation feature position, and a selective L2 distance for efficient network compression.
+ReviewKD [7] leveraged cross-layer feature aggregation to guide the learning of a student model.
 SimKD [6] reused the teacher’s classifier and introduced an additional projector to align student features.
-CAT-KD [17] enhances knowledge distillation’s interpretability by distilling class activation maps.
-This approach empowers the student model to identify discriminative regions, improving its performance.
-Feature-based approaches usually outperform logit-based approaches at the expense of additional computational and storage overhead.
+CAT-KD [17] enhanced the interpretability of knowledge distillation approaches by distilling class activation maps to highlight discriminative regions.
+While feature-based distillation generally outperforms logit-based methods in accuracy, it incurs significant computational overhead due to intermediate feature storage and alignment.
+Existing contrastive knowledge distillation methods typically align teacher-student feature representations by enforcing intra-class similarity and inter-class separation through contrastive learning frameworks.
+For example, CRD [72] treats teacher features as anchors and leverages large-scale negative samples to push student features away from dissimilar classes.
+CRCD [94], which optimizes complementary relationships to reduce dependency on negatives.
+While these approaches demonstrate theoretical promise, there are several limitations, including heavy reliance on large-scale negative sample pairs, compromised scalability due to memory-intensive operations in dense prediction tasks, and insufficient consideration of the fixed teacher model with gradient conflicts.
+To handle these limitations, we propose a sample-wise contrastive learning framework that simultaneously preserves intra-sample feature similarity and cross-sample semantics with the optimized contrastive triplets, effectively addressing the memory constraints.
 II-B Contrastive Learning
 Contrastive learning stands out as a powerful and popular branch of self-supervised learning, and plenty of efforts have been devoted to contrastive learning [3, 23].
-We merely review relevant works concerned with NCE and InfoNCE.
+We merely review relevant works that are concerned with noise-contrastive estimation (NCE) and InfoNCE.
 Gutmann et al.
-[18] firstly proposed the theory of noise-contrastive estimation, in which the model learns to discriminate between samples from the actual data distribution and those from a noise distribution, thereby revealing distinct features of the data.
+[18] firstly proposed the theory of noise-contrastive estimation, in which the model learns to discriminate between samples from the actual data distribution and those from a noise distribution, revealing distinct features of the data.
 In noise-contrastive estimation, increasing the volume of negative samples can largely reduce reliance on the quality of the noise distribution.
 Sohn et al.
-[64] became the first to introduce the multi-class N-pair loss for deep metric learning.
-This approach eliminates the necessity to recalculate embedding representations in every iteration by allowing the reuse of embedding representations from negative samples.
-InstDisc [78] incorporated a memory bank mechanism that retains instance features from the preceding update step, facilitating negative sampling through integration with this non-parametric memory bank.
-CPC [54] popularized the InfoNCE loss for contrastive representation learning, which enhances similarities between a sample and its corresponding positive samples while decreasing similarities with negative samples.
+[66] extended this framework with the multi-class N-pair loss, enabling efficient reuse of negative embeddings across iterations.
+Subsequent works optimized negative sampling and scalability.
+For example, InstDisc [80] introduced a memory bank to store historical instance features, enabling stable negative sampling without recomputation.
+CPC [55] popularized the InfoNCE loss by formalizing contrastive learning as maximizing mutual information between positive pairs.
 Likewise, CMC [71] leveraged a contrastive loss function to optimize the mutual information across different view representations of the same sample.
-MoCo [21] implemented a moving average of model parameters to update a dynamic dictionary using a queue.
-SimCLR [8] employed data augmentation to generate varied inputs of the same image and employed a contrastive loss to enhance the similarity between representations with different augmentations.
+MoCo [21] decoupled batch size from negative sample size via a momentum-updated queue and conducted large-scale negative sampling with limited memory.
+SimCLR [8] demonstrated that strong data augmentation paired with a nonlinear projection head achieves state-of-the-art results, albeit with high computational demands.
 CLIP [59] proposed a large-scale multimodal model pre-trained through the joint training of a text encoder and an image encoder with the InfoNCE loss.
-Previous methods have witnessed the success of contrastive learning.
-However, most contrastive learning approaches require substantial batch sizes of negative samples to optimize training.
-In this paper, we harness the potential of contrastive learning and design sample-wise triples for efficient and effective training.
+While these methods highlight the efficacy of contrastive learning, they rely heavily on large batches of negative samples to ensure the effective separation of representations.
+This dependency creates computational bottlenecks, particularly in dense prediction tasks.
+In this paper, we adapt contrastive learning to knowledge distillation by rethinking negative sampling efficiency.
+Instead of relying on massive negative batches, we propose sample-wise triples that leverage intra-class and inter-class relationships.
+Our approach maintains the structural benefits of contrastive learning while reducing memory overhead, making it scalable to complex tasks like segmentation and detection.
 III Approach
@@
 III-A Notations
-Knowledge distillation tasks aim to transfer knowledge from the teacher model $\mathcal{T}$ to the student model $\mathcal{S}$.
-To be specific, given inputs $\mathcal{X}=[\boldsymbol{x}_{0},\boldsymbol{x}_{1},\ldots,\boldsymbol{x}_{n-1}]$, we calculate teacher logits $\boldsymbol{T}=\mathcal{T}\left(\mathcal{X};\theta_{\mathcal{T}}\right)$ and student logits $\boldsymbol{S}=\mathcal{S}\left(\mathcal{X};\theta_{\mathcal{S}}\right)$, where $\boldsymbol{T}=[\boldsymbol{t}_{0},\boldsymbol{t}_{1},\ldots,\boldsymbol{t}_{n-1}]\in\mathbb{R}^{c\times n}$ and $\boldsymbol{S}=[\boldsymbol{s}_{0},\boldsymbol{s}_{1},\ldots,\boldsymbol{s}_{n-1}]\in\mathbb{R}^{c\times n}$.
-Here, $\theta_{\mathcal{T}}$ and $\theta_{\mathcal{S}}$ are model parameters.
-$\mathbf{t}_{i}$ and $\mathbf{s}_{i}$ represent the teacher and student logits for the $i$-th sample in which $i\in\mathcal{D}$, $\mathcal{D}=[1,2,\ldots,n-1]$.
-$n$ denotes the sample count and $c$ signifies the logit dimension that corresponds to the number of classes.
-Formally, the objective of knowledge distillation approaches can be written as:
+Knowledge distillation is a technique used to transfer knowledge from a trained teacher model $\mathcal{T}$ to a more compact student model $\mathcal{S}$.
+Here are some notations, given a batch of input samples $\mathcal{X}=[\boldsymbol{x}_{0},\boldsymbol{x}_{1},\ldots,\boldsymbol{x}_{n-1}]$, we calculate teacher logits $\boldsymbol{T}=\mathcal{T}\left(\mathcal{X};\theta_{\mathcal{T}}\right)$ and student logits $\boldsymbol{S}=\mathcal{S}\left(\mathcal{X};\theta_{\mathcal{S}}\right)$, where $\boldsymbol{T}=[\boldsymbol{t}_{0},\boldsymbol{t}_{1},\ldots,\boldsymbol{t}_{n-1}]\in\mathbb{R}^{c\times n}$ and $\boldsymbol{S}=[\boldsymbol{s}_{0},\boldsymbol{s}_{1},\ldots,\boldsymbol{s}_{n-1}]\in\mathbb{R}^{c\times n}$.
+$\theta_{\mathcal{T}}$ and $\theta_{\mathcal{S}}$ are model parameters.
+$\mathbf{t}_{i}$ and $\mathbf{s}_{i}$ represent the teacher and student the logits for the $i$-th sample in which $i\in\mathcal{D}$, $\mathcal{D}=[1,2,\ldots,n-1]$.
+$n$ denotes the sample count, and $c$ is the number of classes that corresponds to the logit dimension.
+The student model $\mathcal{S}$ learns to mimic the teacher’s output.
+A typical distillation loss combines task loss and distillation loss.
+Formally, the objective of knowledge distillation approaches can be written as,
 $\mathcal{L}=\mathcal{L}_{\text{Task}}+\alpha\mathcal{L}_{\text{KD}}.$ (1)
-$\mathcal{L}_{\text{Task}}$ represents the standard cross-entropy loss for image classification, and $\alpha$ balances the classification loss $\mathcal{L}_{\text{Task}}$ and the distillation loss $\mathcal{L}_{\text{KD}}$.
-In what follows, we elaborate on the proposed knowledge distillation loss $\mathcal{L}_{\text{KD}}$.
+$\mathcal{L}_{\text{Task}}$ is the standard cross-entropy loss for image classification, while $\alpha$ balances the classification loss $\mathcal{L}_{\text{Task}}$ and the distillation loss $\mathcal{L}_{\text{KD}}$.
+In what follows, we will detail the knowledge distillation loss $\mathcal{L}_{\text{KD}}$.
 III-B Contrastive Knowledge Distillation
-Generally, traditional knowledge distillation methods highlight class-level contrastive information and capture intra-class and inter-class relationships.
-Differently, our method focuses on sample-level representations.
-To be specific, the proposed approach aligns a student logit $\mathbf{s}_{i}$ with its corresponding teacher logit $\mathbf{t}_{i}$ while pushing it apart from other negative logits $\mathbf{s}_{j}$.
-To achieve this goal, we endeavor to distill not only intra-sample numerical similarities but also inter-sample semantic structure.
-Fig.
-2 presents the overall architecture of CKD.
-We can see that our method conducts intra-sample distillation and inter-sample distillation.
-Afterward, we describe these constraints in detail.
+Traditional knowledge distillation methods emphasize class-level contrastive learning by leveraging intra-class and inter-class relationships.
+In contrast, our approach adopts sample-level distillation.
+To be specific, for each sample, we enforce alignment between the student logit $\mathbf{s}_{i}$ with its corresponding teacher logit $\mathbf{t}_{i}$, while simultaneously distancing $\mathbf{s}_{i}$ from negative student logits $\mathbf{s}_{j}\;(j\neq i)$.
+This dual mechanism allows our method to distill not only numerical similarities in individual samples but also semantic structural relationships across different samples.
+As illustrated in Fig.
+2, our contrastive knowledge distillation framework integrates two complementary components.
+The intra-sample distillation directly aligns paired student-teacher logits for each sample, and the inter-sample distillation employs contrastive separation between the target logit $\mathbf{s}_{i}$ and irrelevant logits $\mathbf{s}_{j}$.
+In the following sections, we elaborate on the formulation of these constraints.
 III-B1 Intra-Sample Distillation
-Class-wise methods usually minimize intra-class distances among samples for knowledge transfer.
-By adhering to these methods, the proposed approach maximizes similarities between teacher and student logits for each sample while leveraging sample-wise information.
-Here, we utilize the intra-sample loss, denoted as $\mathcal{L}_{\text{intra}}$, to optimize parameters of the student model $\mathcal{S}$, where $d(\cdot,\cdot)$ represents a distance metric.
-Hence, we derive:
+Unlike class-wise methods that minimize intra-class distances for knowledge transfer, our method explicitly leverages sample-wise information to enhance the similarity between teacher and student logits for each individual sample.
+Here, we introduce the intra-sample loss, denoted as $\mathcal{L}_{\text{intra}}$, to optimize the parameters of the student model $\mathcal{S}$ by minimizing the distance $d(\cdot,\cdot)$ between aligned teacher-student logit pairs.
+Formally, the loss is defined as,
 $\mathcal{L}_{\text{intra}}$ $=\frac{1}{n}\sum_{i=0}^{n}d\left(\boldsymbol{t}_{i},\boldsymbol{s}_{i}\right)$ (2)
 $=\underset{\boldsymbol{x}_{i}\sim\mathcal{X}}{\mathbb{E}}{d\left(\boldsymbol{t}_{i},\boldsymbol{s}_{i}\right)}.$
-For $\boldsymbol{x}_{i}\in\mathcal{X}$, we obtain $\boldsymbol{t}_{i}=\mathcal{T}\left(\boldsymbol{x}_{i};\theta_{\mathcal{T}}\right)$ and $\boldsymbol{s}_{i}=\mathcal{S}\left(\boldsymbol{x}_{i};\theta_{\mathcal{S}}\right)$, where $\boldsymbol{x}_{i}\sim\mathcal{X}$ denotes a data point $\boldsymbol{x}_{i}$ sampled from $\mathcal{X}$.
-The loss function $\mathcal{L}_{\text{intra}}$, defined in Eq.
-2, forces the student logit $\boldsymbol{s}_{i}$ to approximate its analytical solution $\boldsymbol{t}_{i}$.
-Previous studies [59, 55] have shown that continuously minimizing the gap between teacher and student representations can lead to overfitting and reduced generalization to unseen data.
-This is evident in the significant discrepancy between training and validation losses, especially when relying solely on intra-sample similarities.
+For each input sample $\boldsymbol{x}_{i}\in\mathcal{X}$, the teacher and student models generate logits $\boldsymbol{t}_{i}=\mathcal{T}\left(\boldsymbol{x}_{i};\theta_{\mathcal{T}}\right)$ and $\boldsymbol{s}_{i}=\mathcal{S}\left(\boldsymbol{x}_{i};\theta_{\mathcal{S}}\right)$, respectively.
+$\boldsymbol{x}_{i}\sim\mathcal{X}$ means that a data point $\boldsymbol{x}_{i}$ sampled from the data space $\mathcal{X}$.
+In Eq.
+2, the intra-sample loss $\mathcal{L}_{\text{intra}}$ enforces alignment between the student logit $\boldsymbol{s}_{i}$ and its analytical solution $\boldsymbol{t}_{i}$.
+While such alignment improves local similarities, earlier research [61, 56] reveals that exclusive reliance on intra-sample similarity risks overfitting as excessive minimization of the teacher-student gap may degrade generalization to unseen data.
+A phenomenon particularly evident in scenarios dominated by sample-specific optimization.
 (a) $d(\mathbf{t}_{i},\mathbf{s}_{i})<d(\mathbf{t}_{j},\mathbf{s}_{i})$
@@
 The blue line indicates that $\boldsymbol{s}_{i}$ approaches $\boldsymbol{t}_{i}$, and the red line represents that $\boldsymbol{s}_{i}$ is far from $\boldsymbol{s}_{j}$ or $\boldsymbol{t}_{j}$.
+Figure 5: Pseudo code of CKD in a Numpy-like style.
 Fig.
-3 provides a detailed analysis and visualization of the intra-sample constraint on logits $\boldsymbol{t}_{i}$, $\boldsymbol{t}_{j}$, and $\boldsymbol{s}_{i}$.
-The error between the $i$-th teacher and student logits is represented as $\epsilon_{i}=\boldsymbol{t}_{i}-\boldsymbol{s}_{i}$, where $\boldsymbol{t}_{i}\neq\boldsymbol{t}_{j}$.
-The objective $\mathcal{L}_{\text{intra}}$ cannot guarantee $\forall I,d(\boldsymbol{t}_{i},\boldsymbol{s}_{i})>d(\boldsymbol{t}_{j},\boldsymbol{s}_{i})$.
-Therefore, there exist errors $\epsilon_{i}$, such that ${\exists}i,d(\boldsymbol{t}_{i},\boldsymbol{s}_{i})>d(\boldsymbol{t}_{j},\boldsymbol{s}_{i})$.
-This inequality suggests that the student logit $\boldsymbol{s}_{i}$ may be closer to a teacher logit $\boldsymbol{t}_{j}$ than its corresponding teacher logit $\boldsymbol{t}_{i}$.
-Formally, we assume that there are hyperplanes $\mathcal{H}^{(k)},k\in\mathbb{N}$ related to $\boldsymbol{s}_{i}^{(k)}\in\mathcal{H}^{(k)},i\in\mathcal{D}$ with the same loss value, that is:
-$\forall k,h\;\;\sum_{i\in\mathcal{D}}d\left(\boldsymbol{t}_{i},\boldsymbol{s}_{i}^{(k)}\right)=\sum_{i\in\mathcal{D}}d\left(\boldsymbol{t}_{i},\boldsymbol{s}_{i}^{(h)}\right).$ (3)
-The hyperplane $\mathcal{H}^{(k)}$ is defined as:
-$\mathcal{H}^{(k)}:\mathbf{w}^{(k)}\boldsymbol{s}^{(k)}_{i}+b^{(k)}=0,\;i\in\mathcal{D},$ (4)
-where $\boldsymbol{w}\in\mathbb{R}^{c}$ is a normal vector and $b$ is a bias.
-Consider a two-dimensional plane $y=x$ and a teacher logit $\left[0.5,0.5\right]$.
-Despite student logits $\left[0.4,0.4\right]$ and $\left[0.6,0.6\right]$ yielding equal losses under the mean square error metric, they convey different meanings.
-Specifically, the logit $\left[0.4,0.4\right]$ lies below the decision plane, while $\left[0.6,0.6\right]$ lies above it.
-This highlights potential overfitting in intra-sample distillation.
-To address this, additional constraints are necessary.
+3 visualizes the limitations of the intra-sample constraint by analyzing teacher logits $\boldsymbol{t}_{i}$ and student logits $\boldsymbol{t}_{j}$, where $\boldsymbol{t}_{i}\neq\boldsymbol{t}_{j}$.
+We define the alignment error for the $i$-th sample as $\epsilon_{i}=\boldsymbol{t}_{i}-\boldsymbol{s}_{i}$.
+While $\mathcal{L}_{\text{intra}}$ minimizes $d(\boldsymbol{t}_{i},\boldsymbol{s}_{i})$, it does not guarantee the inequality $\forall i,d(\boldsymbol{t}_{i},\boldsymbol{s}_{i})>d(\boldsymbol{t}_{j},\boldsymbol{s}_{i})$.
+Consequently, there may exist cases where $\epsilon_{i}$, such that ${\exists}i,d(\boldsymbol{t}_{i},\boldsymbol{s}_{i})>d(\boldsymbol{t}_{j},\boldsymbol{s}_{i})$.
+This inequality implies that the student logit $\boldsymbol{s}_{i}$ becomes closer to an incorrect teacher logit $\boldsymbol{t}_{j}$ than to its assigned target $\boldsymbol{t}_{i}$.
+Formally, we consider a family of hyperplanes $\mathcal{H}^{(k)},k\in\mathbb{N}$ in the student logit space, where each $\boldsymbol{s}_{i}^{(k)}$ under $\mathcal{H}^{(k)}$ satisfies $\boldsymbol{s}_{i}^{(k)}\in\mathcal{H}^{(k)},i\in\mathcal{D}$.
+These hyperplanes share the same aggregated loss value across all samples as,
+$\forall k,h\in\mathbb{N}\;\;\sum_{i\in\mathcal{D}}d\left(\boldsymbol{t}_{i},\boldsymbol{s}_{i}^{(k)}\right)=\sum_{i\in\mathcal{D}}d\left(\boldsymbol{t}_{i},\boldsymbol{s}_{i}^{(h)}\right).$ (3)
+Each hyperplane $\mathcal{H}^{(k)}$ is parameterized as,
+$\mathcal{H}^{(k)}:\boldsymbol{w}^{(k)}\boldsymbol{s}^{(k)}_{i}+b^{(k)}=0,\;i\in\mathcal{D},$ (4)
+where $\boldsymbol{w}^{(k)}\in\mathbb{R}^{c}$ is the normal vector and $b$ is the bias term.
+Consider a binary classification scenario with a decision boundary defined by $y=x$.
+Suppose a teacher logs $\mathbf{t}_{i}=\left[0.5,0.5\right]$ corresponds to a sample near this boundary.
+Two student logits $\left[0.4,0.4\right]$ and $\left[0.6,0.6\right]$ yield identical intra-sample losses under mean squared error.
+However, their geometric positions relative to the decision boundary differ critically.
+$\left[0.4,0.4\right]$ falls below the boundary, implying misclassification.
+$\left[0.6,0.6\right]$ lies above the boundary, preserving correct classification.
+These loss values expose a key limitation of intra-sample distillation that $\mathcal{L}_{intra}$ may preserve numerical similarities while degrading task-specific correctness.
+To mitigate such overfitting risks, we introduce supplementary constraints that enforce semantic consistency beyond intra-sample alignment.
 III-B2 Inter-Sample Distillation
-Knowledge distillation should guarantee numerical similarity and semantic consistency between teacher logits and corresponding student logits [70, 4].
-However, intra-sample alignment only emphasizes numerical similarity to align target logits.
-We introduce another inter-sample semantic alignment to simultaneously separate the logits of different samples.
-The inter-sample constraint $\mathcal{L}_{\text{inter}}$ is defined as:
+Knowledge distillation should guarantee both numerical similarity and semantic consistency between teacher and student logits [72, 4].
+However, conventional intra-sample distillation focuses on the former by directly aligning paired logits.
+To address this limitation, we introduce inter-sample semantic distillation, which explicitly enforces structural relationships across samples.
+The inter-sample constraint $\mathcal{L}_{\text{inter}}$ operates complementarily to $\mathcal{L}_{\text{intra}}$ through contrastive separation by pushing student logits of distinct samples apart in the embedding space.
+Formally, $\mathcal{L}_{\text{inter}}$ is defined as,
 $\mathcal{L}_{\text{inter}}$ $=-\frac{1}{n(n-1)}\sum_{i=0}^{n}\sum_{j\neq i}^{n}\left[d\left(\boldsymbol{s}_{i},\boldsymbol{s}_{j}\right)+d\left(\boldsymbol{s}_{i},\boldsymbol{t}_{j}\right)\right]$ (5)
 $=-\underset{\begin{subarray}{c}{\boldsymbol{x}_{i}\sim\mathcal{X}},{\boldsymbol{x}_{j}\sim\mathcal{X}}\end{subarray}}{\mathbb{E}}{d\left(\boldsymbol{s}_{i},\boldsymbol{s}_{j}\right)}-\underset{\begin{subarray}{c}{\boldsymbol{x}_{i}\sim\mathcal{X}},{\boldsymbol{x}_{j}\sim\mathcal{X}}\end{subarray}}{\mathbb{E}}{d\left(\boldsymbol{s}_{i},\boldsymbol{t}_{j}\right)}.$
-where $\boldsymbol{x}_{i}\neq\boldsymbol{x}_{j}$.
-The objective $\mathcal{L}_{\text{inter}}$ is to make $\boldsymbol{s}_{i}$ far from both $\boldsymbol{s}_{j}$ and $\boldsymbol{t}_{j}$.
+The objective loss $\mathcal{L}_{\text{inter}}$ enforces dual separation that it pushes the student logit $\boldsymbol{s}_{i}$ away from other student logits $\boldsymbol{s}_{j}$ and their corresponding teacher logits $\boldsymbol{t}_{j}$.
 As shown in Fig.
-3(b), we can alleviate the misalignment of $\boldsymbol{s}_{i}$ and $\boldsymbol{t}_{j}$ with Eq.
-5 by maximizing their distance simultaneously.
-Generally, inter-sample pairs include $\left(\boldsymbol{s}_{i},\boldsymbol{s}_{j}\right)$ and $\left(\boldsymbol{s}_{i},\boldsymbol{t}_{j}\right)$.
-We consider intra-sample pairs and can obtain triples $\left(\boldsymbol{s}_{i},\boldsymbol{t}_{i},\boldsymbol{s}_{j}\right)$ and $\left(\boldsymbol{s}_{i},\boldsymbol{t}_{i},\boldsymbol{t}_{j}\right)$.
+3(b), this constraint resolves cross-sample misalignment by preventing $\boldsymbol{s}_{i}$ from collapsing towards irrelevant teacher logits $\boldsymbol{t}_{j}$ and inter-sample redundancy by ensuring discriminative separation between distinct student logits $\boldsymbol{s}_{i}$ and $\boldsymbol{s}_{j}$.
+Eq.
+5 achieves this through simultaneous distance maximization between $\left(\boldsymbol{s}_{i},\boldsymbol{s}_{j}\right)$ and $\left(\boldsymbol{s}_{i},\boldsymbol{t}_{j}\right)$.
+Generally, for each student logit $\boldsymbol{s}_{i}$, we construct inter-sample pairs include student-student pairs $\left(\boldsymbol{s}_{i},\boldsymbol{s}_{j}\right)$ or student-teacher pairs $\left(\boldsymbol{s}_{i},\boldsymbol{t}_{j}\right)$.
+We combine the intra-sample pair $\left(\boldsymbol{s}_{i},\boldsymbol{t}_{i}\right)$ and obtain two types of triples $\left(\boldsymbol{s}_{i},\boldsymbol{t}_{i},\boldsymbol{s}_{j}\right)$ and $\left(\boldsymbol{s}_{i},\boldsymbol{t}_{i},\boldsymbol{t}_{j}\right)$.
 Fig.
 4 visualizes these two triple categories.
-Previous knowledge distillation approaches [70, 4] use both triple types for training.
-However, unlike these methods [64, 21, 8] in which each triple element is trainable, the teacher model remains fixed, with its output $\boldsymbol{t}_{i}$ unchanged.
-It may be questionable to simply apply these two triples in knowledge distillation methods.
-To address this issue, we only use triples $\left(\boldsymbol{s}_{i},\boldsymbol{t}_{i},\boldsymbol{s}_{j}\right)$ during training.
+While previous triple-based methods [72, 4] leverage both triple types by updating all elements [66, 21, 8], our framework differs as the teacher model remains fixed with its outputs $\boldsymbol{t}_{i}$ unchanged during distillation.
+This constraint introduces a critical challenge that the direct adoption of traditional triplet losses designed for fully trainable models may lead to suboptimal alignment when applied to static teacher outputs.
+Our method exclusively adopts $\left(\boldsymbol{s}_{i},\boldsymbol{t}_{i},\boldsymbol{s}_{j}\right)$ for distillation.
 Fig.
-4 briefly illustrates the rationale behind this selection.
+4 briefly illustrates the underlying reasoning behind this selection.
 In Fig.
-4(a), $\boldsymbol{s}_{i}$ approaches $\boldsymbol{t}_{i}$ for intra-sample alignment while distancing itself from $\boldsymbol{s}_{j}$ for inter-sample alignment.
-Even with a gradient conflict when $\theta_{i}>\pi/2$, we can resolve it by adjusting $\boldsymbol{s}_{j}$.
+4(a), the student logit $\boldsymbol{s}_{i}$ approaches $\boldsymbol{t}_{i}$ for intra-sample alignment while distancing from $\boldsymbol{s}_{j}$ for inter-sample separation.
+While gradient conflicts may emerge when the angle $\theta_{i}>\pi/2$, these can be mitigated by adjusting the trainable $\boldsymbol{s}_{j}$ during optimization.
 However, in Fig.
-4(b), the conflict remains unresolved due to the fixed $\boldsymbol{t}_{j}$.
-Additionally, Fig.
-4(c) depicts the worst-case scenario where the gradient conflict is at its maximum.
+4(b), the conflict remains unresolved due to the immutable nature of $\boldsymbol{t}_{j}$.
+In extreme casesFig.
+4(c), opposing gradients from $\boldsymbol{t}_{i}$ and $\boldsymbol{t}_{j}$ destabilize training, as neither term can be adaptively regularized.
 From Fig.
-4, we can conclude that inter-samples are crucial for model updates and the triple $\left(\boldsymbol{s}_{i},\boldsymbol{t}_{i},\boldsymbol{t}_{j}\right)$ may cause gradient conflicts.
-Thus, our method merely adopt the triple $\left(\boldsymbol{s}_{i},\boldsymbol{t}_{i},\boldsymbol{s}_{j}\right)$, and the inter-sample objective is formed as:
+4, we conclude that inter-sample interactions must involve trainable student logits to ensure stable optimization and that the triple $\left(\boldsymbol{s}_{i},\boldsymbol{t}_{i},\boldsymbol{t}_{j}\right)$ may induce gradient conflicts.
+Consequently, our method exclusively adopts $\left(\boldsymbol{s}_{i},\boldsymbol{t}_{i},\boldsymbol{s}_{j}\right)$, and therefore formulate the inter-sample objective as the following equation,
 $\mathcal{L}_{\text{inter}}$ $=-\frac{1}{n(n-1)}\sum_{i=0}^{n}\sum_{j\neq i}^{n}d\left(\boldsymbol{s}_{i},\boldsymbol{s}_{j}\right)$ (6)
 $=-\underset{\begin{subarray}{c}{\boldsymbol{x}_{i}\sim\mathcal{X}},{\boldsymbol{x}_{j}\sim\mathcal{X}}\end{subarray}}{\mathbb{E}}{d\left(\boldsymbol{s}_{i},\boldsymbol{s}_{j}\right)}.$
-Finally, we combine Eq. 2 and Eq. 6 with a tradeoff parameter $\beta$ to obtain the following loss function:
+Finally, we unify the intra-sample alignment constraint Eq. 2 and inter-sample separation constraint Eq. 6 through a weighted summation by using a balancing parameter $\beta$ as,
 $\mathcal{L}_{\text{KD}}=\mathcal{L}_{\text{intra}}+\beta\mathcal{L}_{\text{inter}}.$ (7)
-Eq.
-7 is to maximize similarities between teacher and student logits for the same sample while minimizing correlations across different samples.
-By pulling positive pairs closer and pushing negative pairs apart, the intra-sample and inter-sample formulations enable the alignment of teacher and student logits, ensuring they share similar semantic information for similarity and dissimilarity.
-Figure 5: Pseudo code of CKD in a Numpy-like style.
+The objective Eq.
+7 achieves intra-sample maximization by enhancing the similarity between teacher and student logits for the same sample and inter-sample minimization by suppressing correlations between the logits of distinct samples.
+This is realized through contrastive learning with positive pairs pulled closer via $\mathcal{L}_{intra}$ and negative pairs pushed apart via $\mathcal{L}_{inter}$.
 III-B3 Contrastive Formulation
-The proposed structure-preserving alignment shares a similar objective with contrastive learning [54, 79].
-We seek to transform our objective into a sample-wise contrastive learning formulation.
-Specifically, $\mathcal{L}_{\text{KD}}$ can be expressed as:
+Notably, the proposed structure-preserving distillation framework exhibits an objective akin to contrastive learning [55, 81].
+This connection motivates us to reformulate our objective into a sample-wise contrastive learning framework.
+Formally, $\mathcal{L}_{\text{KD}}$ can be expressed as,
 $\begin{split}\mathcal{L}_{\text{KD}}&=\underset{\boldsymbol{x}_{i}\sim\mathcal{X}}{\mathbb{E}}{d\left(\boldsymbol{t}_{i},\boldsymbol{s}_{i}\right)}-\beta\underset{\begin{subarray}{c}{\boldsymbol{x}_{i}\sim\mathcal{X}}\\ {\boldsymbol{x}_{j}\sim\mathcal{X}}\end{subarray}}{\mathbb{E}}{d\left(\boldsymbol{s}_{i},\boldsymbol{s}_{j}\right)}.\end{split}$ (8)
-For simplicity, we utilize the distance metric $f(\boldsymbol{u},\boldsymbol{v})$ to assess the similarity between $\boldsymbol{u}$ and $\boldsymbol{v}$, and set $f(\boldsymbol{u},\boldsymbol{v})=-d(\boldsymbol{u},\boldsymbol{v})$.
-The loss function can be derived as:
-$\mathcal{L}_{\text{KD}}$ $=\underset{\boldsymbol{x}_{i}\sim\mathcal{X}}{\mathbb{E}}\left[-(f\left(\boldsymbol{t}_{i},\boldsymbol{s}_{i}\right)-\beta\underset{\boldsymbol{x}_{j}\sim\mathcal{X}}{\mathbb{E}}f\left(\boldsymbol{s}_{i},\boldsymbol{s}_{j}\right))\right]$ (9)
-$=\underset{\boldsymbol{x}_{i}\sim\mathcal{X}}{\mathbb{E}}\ln\left[{\frac{\underset{\boldsymbol{x}_{j}\sim\mathcal{X}}{\mathbb{E}}e^{\beta f\left(\boldsymbol{s}_{i},\boldsymbol{s}_{j}\right)}}{e^{f\left(\boldsymbol{t}_{i},\boldsymbol{s}_{i}\right)}}}+\frac{e^{f\left(\boldsymbol{t}_{i},\boldsymbol{s}_{i}\right)}}{e^{f\left(\boldsymbol{t}_{i},\boldsymbol{s}_{i}\right)}}-1\right]$ (10)
+For simplicity, we employ the distance metric $f(\boldsymbol{u},\boldsymbol{v})$ to quantify the similarity between $\boldsymbol{u}$ and $\boldsymbol{v}$, defining $f(\boldsymbol{u},\boldsymbol{v})=-d(\boldsymbol{u},\boldsymbol{v})$.
+The loss function can be derived as,
+$\mathcal{L}_{\text{KD}}$ $=\underset{\boldsymbol{x}_{i}\sim\mathcal{X}}{\mathbb{E}}\left[-(f\left(\boldsymbol{t}_{i},\boldsymbol{s}_{i}\right)-\beta\underset{\boldsymbol{x}_{j}\sim\mathcal{X}}{\mathbb{E}}f\left(\boldsymbol{s}_{i},\boldsymbol{s}_{j}\right))\right].$ (9)
+where higher values indicate greater similarity.
+Employing Eq.
+9 directly for optimization may result in gradient conflicts or diminish training efficiency.
+To rewrite it in the format of contrastive loss, we incorporate the natural logarithm for derivation as,
+$\mathcal{L}_{\text{KD}}$ $=\underset{\boldsymbol{x}_{i}\sim\mathcal{X}}{\mathbb{E}}\ln\left[{\frac{\underset{\boldsymbol{x}_{j}\sim\mathcal{X}}{\mathbb{E}}e^{\beta f\left(\boldsymbol{s}_{i},\boldsymbol{s}_{j}\right)}}{e^{f\left(\boldsymbol{t}_{i},\boldsymbol{s}_{i}\right)}}}+\frac{e^{f\left(\boldsymbol{t}_{i},\boldsymbol{s}_{i}\right)}}{e^{f\left(\boldsymbol{t}_{i},\boldsymbol{s}_{i}\right)}}-1\right]$ (10)
 $\simeq\underset{\boldsymbol{x}_{i}\sim\mathcal{X}}{\mathbb{E}}\left[-\ln{\frac{e^{f\left(\boldsymbol{t}_{i},\boldsymbol{s}_{i}\right)}}{e^{f\left(\boldsymbol{t}_{i},\boldsymbol{s}_{i}\right)}+\underset{\boldsymbol{x}_{j}\sim\mathcal{X}}{\mathbb{E}}e^{\beta f\left(\boldsymbol{s}_{i},\boldsymbol{s}_{j}\right)}}}\right]$ (11)
 $\simeq\underset{\boldsymbol{x}_{i}\sim\mathcal{X}}{\mathbb{E}}\left[-\ln{\frac{e^{f\left(\boldsymbol{t}_{i},\boldsymbol{s}_{i}\right)/\tau}}{e^{f\left(\boldsymbol{t}_{i},\boldsymbol{s}_{i}\right)/\tau}+\underset{\boldsymbol{x}_{j}\sim\mathcal{X}}{\mathbb{E}}e^{\beta f\left(\boldsymbol{s}_{i},\boldsymbol{s}_{j}\right)/\tau}}}\right].$ (12)
-The objective can be viewed as a sample-wise contrastive loss.
 Here, the teacher logit $\boldsymbol{t}_{i}$ and its corresponding student logit $\boldsymbol{s}_{i}$ form a positive pair, while $\boldsymbol{s}_{i}$ and $\boldsymbol{s}_{j}$ form a negative pair.
 Next, we will explore the benefits of this contrastive formulation.
-(a) Traditional class-wise contrastive
-(b) Typical class-wise contrastive: CRD
-(c) Proposed sample-wise contrastive
+(a) Traditional class-wise contrast
+(b) Typical class-wise contrast: CRD
+(c) Proposed sample-wise contrast
 Figure 6: Visualization of two categories of contrastive learning: Class-wise vs.
@@
 As shown in Eq.
-9, to calculate the expected value $f(\boldsymbol{s}_{i},\boldsymbol{s}_{j})$ for each $\boldsymbol{x}_{i}$, we sample negative pairs from the data space $\boldsymbol{x}_{j}$.
-Following vanilla KD [25], we introduce a temperature parameter $\tau$ to soften the probability distribution, satisfying Eq.
+9, to compute the expected value $f(\boldsymbol{s}_{i},\boldsymbol{s}_{j})$ for each $\boldsymbol{x}_{i}$, we sample negative pairs from the data space $\mathcal{X}$.
+Following vanilla KD [25], we introduce a temperature parameter $\tau$ to soften the logit distribution defined in Eq.
 12.
-The parameter $\beta$ balances the importance of positive and negative pairs, where a larger $\beta$ is assigned when negative pairs are important; otherwise, a smaller value is preferable.
-Note that negative sampling can also achieve this trade-off.
-When the negative space is significant, we increase the sampling of negative pairs for more detailed space estimation and weight adjustment.
-Conversely, fewer negative samples are selected when they are less important.
-Contrastive learning enables efficient batch data processing in parallel.
-In our formulation, we construct batches of training data, each containing one positive sample and $B-1$ negative samples, where $B$ is the batch size.
-From Fig.
+The parameter $\beta$ dynamically adjusts the trade-off between intra-sample alignment and inter-sample separation.
+We assign larger $\beta$ values when semantic separation through negative pairs is critical and smaller values otherwise.
+Note that this balance can also be achieved through adaptive negative sampling.
+For example, in scenarios with dominant negative spaces, we increase the number of sampled negative pairs to improve space estimation and weight adjustment.
+When negative pairs are less informative, we reduce their sampling frequency.
+Contrastive learning enables efficient parallel processing of batch data.
+In our formulation, we construct training batches of data samples, each containing one positive sample and $B-1$ negative samples, where $B$ denotes the batch size.
+As shown in Fig.
 2, the number of negative samples depends on $B$.
-Therefore, by varying $B$, we can sample different numbers of negative samples, balancing intra-sample and inter-sample alignment.
-Additionally, we substitute $f\left(\boldsymbol{t}_{i},\boldsymbol{s}_{j}\right)$ for $f\left(\boldsymbol{s}_{i},\boldsymbol{s}_{j}\right)$.
-First, our method primarily aims to align $\boldsymbol{s}_{i}$ and $\boldsymbol{t}_{i}$, such that $\boldsymbol{t}_{i}=\boldsymbol{s}_{i}$.
-This leads to $\boldsymbol{t}_{i}=\boldsymbol{s}_{i}\Rightarrow f\left(\boldsymbol{t}_{i},\boldsymbol{s}_{j}\right)=f\left(\boldsymbol{s}_{i},\boldsymbol{s}_{j}\right)$.
-Then, $f\left(\boldsymbol{s}_{i},\boldsymbol{s}_{j}\right)$ involves two updating variables but $f\left(\boldsymbol{t}_{i},\boldsymbol{s}_{j}\right)$ only has one variable, simplifying the optimization of $\boldsymbol{s}_{j}$.
-Specifically, we treat $\boldsymbol{t}_{i}$ as an anchor logit and separately learn its positive logit $\boldsymbol{s}_{i}$ and negative logit $\boldsymbol{s}_{j}$.
-Most importantly, this setup allows our method to leverage negative samples through the N-pair loss [64] or InfoNCE loss [54].
-Similar to CLIP [58], we sample a batch of $B$ (teacher, student) pairs with $B\times B$ (teacher, student) pairings for efficient training.
+Therefore, adjusting $B$ thus provides explicit control over the intra-inter sample balance.
+Additionally, we replace $f\left(\boldsymbol{t}_{i},\boldsymbol{s}_{j}\right)$ with $f\left(\boldsymbol{s}_{i},\boldsymbol{s}_{j}\right)$.
+This substitution is justified by two key considerations.
+First, our method primarily aligns $\boldsymbol{s}_{i}$ and $\boldsymbol{t}_{i}$, such that $\boldsymbol{t}_{i}=\boldsymbol{s}_{i}$.
+We have $\boldsymbol{t}_{i}=\boldsymbol{s}_{i}\Rightarrow f\left(\boldsymbol{t}_{i},\boldsymbol{s}_{j}\right)=f\left(\boldsymbol{s}_{i},\boldsymbol{s}_{j}\right)$.
+Second, $f\left(\boldsymbol{s}_{i},\boldsymbol{s}_{j}\right)$ involves two trainable variables, whereas $f\left(\boldsymbol{t}_{i},\boldsymbol{s}_{j}\right)$ depends only on $\boldsymbol{s}_{j}$.
+This optimization simplifies gradient propagation through $\boldsymbol{s}_{j}$.
+Specifically, we treat $\boldsymbol{t}_{i}$ as an anchor logit, and its positive logit $\boldsymbol{s}_{i}$ and negative logit $\boldsymbol{s}_{j}$ serve as learnable positive and negative logits, respectively.
+This design naturally connects with the InfoNCE loss [55], enabling batch-efficient training through $B\times B$ (teacher, student) pairings per batch analogous to CLIP [59].
+Here, batch size $B$ controls the number of negative samples, balancing alignment accuracy and computational cost.
 Consequently, we derive the final loss function as,
@@
 Theorem 1.
-Given triples $\left(\boldsymbol{t}_{i},\boldsymbol{s}_{i},\boldsymbol{s}_{j}\right)$ and the loss function $\mathcal{L}_{i}$ that is defined as:
+Given triples $\left(\boldsymbol{t}_{i},\boldsymbol{s}_{i},\boldsymbol{s}_{j}\right)$, the loss function $\mathcal{L}_{i}$ is defined as,
 $\mathcal{L}_{i}=-\ln{\frac{e^{f\left(\boldsymbol{t}_{i},\boldsymbol{s}_{i}\right)/\tau}}{e^{f\left(\boldsymbol{t}_{i},\boldsymbol{s}_{i}\right)/\tau}+{\textstyle\sum\limits^{B}}e^{f\left(\boldsymbol{t}_{i},\boldsymbol{s}_{j}\right)/\tau}}}.$ (15)
-The gradients $\nabla_{\boldsymbol{s}_{i}}\mathcal{L}_{i}$ and $\nabla_{\boldsymbol{s}_{j}}\mathcal{L}_{i}$ are proportional to $g_{i}$ that is formed as:
+The gradients $\nabla_{\boldsymbol{s}_{i}}\mathcal{L}_{i}$ and $\nabla_{\boldsymbol{s}_{j}}\mathcal{L}_{i}$ are proportional to $g_{i}$ that is formed as,
 $g_{i}=1-\frac{e^{f\left(\boldsymbol{t}_{i},\boldsymbol{s}_{i}\right)/\tau}}{e^{f\left(\boldsymbol{t}_{i},\boldsymbol{s}_{i}\right)/\tau}+{\textstyle\sum\limits^{B}}e^{f\left(\boldsymbol{t}_{i},\boldsymbol{s}_{j}\right)/\tau}}.$ (16)
@@
 For class-wise contrastive learning, $\boldsymbol{x}_{i}$ and $\boldsymbol{x}_{j}$ belong to different classes and $f\left(\boldsymbol{t}_{i},\boldsymbol{s}_{i}\right)-f\left(\boldsymbol{t}_{i},\boldsymbol{s}_{j}\right)=\rho_{j}$.
-When we decrease $B$, The gradients $\nabla_{\boldsymbol{s}_{i}}\mathcal{L}_{i}$ also decreases.
+When decreasing $B$, the gradients $\nabla_{\boldsymbol{s}_{i}}\mathcal{L}_{i}$ also decreases.
 Proof.
-From Eq. 16, we can obtain following equations:
+From Eq. 16, we can obtain following equations,
 $g_{i}$ $=1-\frac{e^{f\left(\boldsymbol{t}_{i},\boldsymbol{s}_{i}\right)/\tau}}{e^{f\left(\boldsymbol{t}_{i},\boldsymbol{s}_{i}\right)/\tau}+{\textstyle\sum\limits^{B}}e^{f\left(\boldsymbol{t}_{i},\boldsymbol{s}_{j}\right)/\tau}}$ (17)
@@
 $\simeq 1-\frac{1}{1+B\times e^{-\rho_{j}/\tau}}.$ (19)
-When $B$ decreases, we sample fewer negative samples, resulting in a larger $\rho_{j}$ and a smaller $g_{i}$. Consequently, the gradients $\nabla_{\boldsymbol{s}_{i}}\mathcal{L}_{i}$ decrease. ∎
-From Theorem 2, we can find that when positive samples and negative samples exhibit weak correlations, the value of $f(\boldsymbol{t}_{i},\boldsymbol{s}_{j})$ is minimized, reducing the gradient of backpropagation $\nabla_{\boldsymbol{s}_{i}}\mathcal{L}_{i}$.
-This hampers model parameter updates.
-Therefore, class-wise methods [64, 21] heavily rely on large batch sizes to mitigate this issue and improve learning efficiency [84].
+When the batch size $B$ decreases, fewer negative samples are sampled, leading to an increase in $\rho_{j}$ and a decrease in $g_{i}$.
+This, in turn, results in a reduction of gradients $\nabla_{\boldsymbol{s}_{i}}\mathcal{L}_{i}$.
+∎
+As proven in Theorem 2, when positive and negative samples exhibit weak correlations, the similarity measure $f(\boldsymbol{t}_{i},\boldsymbol{s}_{j})$ approaches its minimum.
+This suppresses the backpropagation gradient magnitude $\nabla_{\boldsymbol{s}_{i}}\mathcal{L}_{i}$, leading to inefficient model parameter updates.
+This limitation forces class-wise contrastive methods [66, 21] to rely on large batch sizes for gradient stabilization by amplifying signal-to-noise ratios through abundant negative samples and for correlation enhancement by mitigating weak correlation scenarios via batch sampling.
+Such strategies, while effective, incur substantial computational costs in contrastive learning [86].
 Input: Images $\mathcal{X}={\{\boldsymbol{x}_{i}\}_{i=1}^{n-1}}$, annotations $\mathcal{Y}={\{y_{i}\}_{i=1}^{n-1}}$, and the teacher model $\mathcal{T}(\mathcal{X};\theta_{\mathcal{T}})$
@@
 Algorithm 1 Training procedure of the proposed CKD
-TABLE I: Classification accuracy (%) on the CIFAR-100 validation set, where teacher and student networks share the same architecture type.
+Theorem 3.
+For sample-wise contrastive learning, $\exists j$ where $\boldsymbol{x}_{i}$ and $\boldsymbol{x}_{j}$ share the same class label, such that $\exists j,\;f\left(\boldsymbol{t}_{i},\boldsymbol{s}_{i}\right)\sim f\left(\boldsymbol{t}_{i},\boldsymbol{s}_{j}\right)$, large batch sizes are not crucial.
+Proof.
+For sample-wise triples where $\exists j,f\left(\boldsymbol{t}_{i},\boldsymbol{s}_{i}\right)\sim f\left(\boldsymbol{t}_{i},\boldsymbol{s}_{j}\right)$, we can obtain $\rho_{j}\sim 0$ and a larger $B\times e^{-\rho_{j}}$ is achieved, when sampling negative samples from the same class.
+Our method eliminates the need for comprehensive negative space sampling with large batch sizes.
+This finding aligns with the theoretical insights established by Robinson et al.[60], who demonstrated that difficult negative samples that are close to the anchor can provide more information and accelerate the gradient update.
+∎
+Some works have revealed that each image is distinctive and can vary significantly from other images within the same semantic class [80, 48].
+It suggests that sample-wise contrastive learning could potentially outperform class-wise approaches by preserving intrinsic structural information in this case.
+Specifically, class-wise methods, which enforce uniform aggregation of intra-class samples toward a centroid, risk distorting the natural semantic structures inherent in the data.
+Other research suggests that negative samples in the same class would harm contrastive learning, as the distance between same-class samples may become excessively large [28].
+Our method addresses this challenge through teacher-guided structural constraints.
+When feature representations $\boldsymbol{s}_{i}$ and $\boldsymbol{s}_{j}$ from the same class move away from each other, their respective convergence towards corresponding teacher anchors $\boldsymbol{t}_{i}$ and $\boldsymbol{t}_{j}$ effectively constrains their relative positions.
+Consequently, this process leads to the alignment between the student’s intra-class semantic structure and that of the teacher.
+Fig.
+6 illustrates the differences between class-wise and sample-wise contrastive paradigms.
+Traditional class-wise approaches like CRD [72] employ memory banks to construct negative pairs through class exclusion, i.e., $\mathbf{x}_{j}\in\mathcal{C}_{k}$ where $k\neq y_{i}$ and $\mathcal{C}_{k}$ is $k$-th category of samples with
+$\mathbb{E}\left[f\left(\boldsymbol{t}_{i},\boldsymbol{s}_{j}\right)\right]\rightarrow 0\quad\forall j\in\mathcal{C}_{k\neq y_{i}}.$ (20)
+The class-wise formulation generates negative pairs from different classes, and these negative samples often exhibit low similarity due to diverse semantic content, forcing gradient enhancement through batch size scaling $g_{i}\propto B$.
+In contrast, the sample-wise formulation may sample negative pairs from the same class as,
+$\mathbb{E}\left[f\left(\boldsymbol{t}_{i},\boldsymbol{s}_{j}\right)\right]\gg 0.$ (21)
+This non-zero similarity naturally amplifies gradient signals without batch size dependency.
+Therefore, our method does not require a large batch size.
+In fact, the proposed method exhibits batch sensitivity.
+Excessive $B$ introduces over-regularization while Insufficient $B$ empirically degrades convergence stability.
+TABLE I: Classification accuracy (%) on the CIFAR-100 validation set.
+”Homogeneous Architectures” and ”Heterogeneous Architectures” represent whether teacher and student networks share the same or different architectures, respectively.
 $\mathrm{\Delta}$ denotes the improved performance compared to vanilla KD.
 All results are the average over five trials.
-Methods Teacher ResNet56 ResNet110 ResNet32$\times$4 WRN-40-2 WRN-40-2 VGG13
-72.34 74.31 79.42 75.61 75.61 74.64
-Student ResNet20 ResNet32 ResNet8$\times$4 WRN-16-2 WRN-40-1 VGG8
-69.06 71.14 72.50 73.26 71.98 70.36
-Features FitNet[59] 69.21 71 06 73.50 73.58 72.24 71.02
-AT[31] 70.55 72.31 73.44 74.08 72.77 71.43
-RKD[55] 69.61 71.82 71.90 73.35 72.22 71.48
-CRD[70] 71.16 73.48 75.51 75.48 74.14 73.94
-OFD[24] 70.98 73.23 74.95 75.24 74.33 73.95
-ReviewKD[7] 71.89 73.89 75.63 76.12 75.09 74.84
-Logits Vanilla KD[25] 70.66 73.08 73.33 74.92 73.54 72.98
-TAKD[51] 70.83 73.37 73.81 75.12 73.78 73.23
-DKD[90] 71.97 74.11 76.32 76.24 74.81 74.68
-CKD 72.12 74.48 76.76 76.28 75.14 74.98
-$\mathrm{\Delta}$ +1.46 +1.40 +3.43 +1.36 +1.60 +2.00
-Theorem 3.
-For sample-wise contrastive learning, $\exists j$, $\boldsymbol{x}_{i}$ and $\boldsymbol{x}_{j}$ are in the same class, such that $\exists j,\;f\left(\boldsymbol{t}_{i},\boldsymbol{s}_{i}\right)\sim f\left(\boldsymbol{t}_{i},\boldsymbol{s}_{j}\right)$, large batch sizes are not crucial.
-Proof.
-For sample-wise triples, $\exists j,f\left(\boldsymbol{t}_{i},\boldsymbol{s}_{i}\right)\sim f\left(\boldsymbol{t}_{i},\boldsymbol{s}_{j}\right)$, we can obtain $\exists j,\rho_{j}\sim 0$ and a larger $B\times e^{-\rho_{j}}$ is achieved, when we sample negative samples that are from the same class.
-Our method is unnecessary for comprehensive negative space sampling with large batch sizes.
-∎
-Fig.
-6 visualizes traditional class-wise and the proposed sample-wise formulations.
-Specifically, CRD [70] is a representative class-wise contrastive learning approach by using a memory bank.
-We can observe that class-wise formulations including CRD generate negative pairs from different classes.
-These negative samples often have low similarity due to varied semantic content, necessitating a larger batch size to increase $g_{i}$.
-In contrast, the sample-wise formulation selects negative samples from the same class, thereby exhibiting stronger relationships.
-Thus, our method does not require a large batch size.
-In fact, a large batch size can hinder our method by influencing intra-sample alignment with a large $\beta$.
-However, a small batch size may also lead to insufficient gradient magnitude.
-Specifically, given the logit space $\boldsymbol{T}\in\mathbb{R}^{c\times n}$, the negative logit space for the $i$-th sample is denoted as $\boldsymbol{T}_{/i}\in\mathbb{R}^{c\times(n-1)}$.
-Assume that there are $c$ classes, each containing $m$ samples.
-We can derive:
-$\text{rank}\left(\boldsymbol{T}_{/i}\right)\leq\min{\left(c,n-1\right)}=c,\;\textit{ s.t. }c\ll n.$ (20)
-Eq.
-20 suggests that in the proposed method, the number of negative pairs sampled should heavily depend on the class number $c$ rather than the sample number $n$.
-Taking category correlations into consideration, the actual number of negative pairs sampled should be even lower than $c$.
-Furthermore, compared to class-wise contrastive learning approaches, our method can be efficiently trained with well-designed triples.
-For classic contrastive learning methods with $n$ samples, a total of $\mathcal{O}(nm)$ triples should be selected for each sample:
-$\binom{m-1}{1}\binom{n-m}{1}=nm-m^{2}-n+m\propto\mathcal{O}(nm)$ (21)
-Eq.
-21 shows that for each sample, we randomly select one sample within the same class including $m-1$ selections, and one sample from a different class, yielding a total of $\mathcal{O}(nm)$ triples.
-As for CRD, we only need to select one sample from different classes because the positive pair is obtained by using the same sample:
-$\binom{1}{1}\binom{n-m}{1}=n-m\propto\mathcal{O}(n).$ (22)
-However, the negative space $\mathcal{O}(n)$ is vast, necessitating the use of a memory bank with a size of $K$.
-As a result, $K$ triples are available for training each sample.
-In contrast, the proposed method requires sampling only $\mathcal{O}(c)$ triples:
-$\binom{1}{1}\binom{c}{1}=c\propto\mathcal{O}(1),$ (23)
-where $c\ll n$ We randomly shuffle data samples and select non-overlapping $c$ samples in order.
-Therefore, our sample-wise formulation exhibits efficient trainability in comparison to the class-wise formulation in contrastive knowledge distillation.
-To sum up, we encapsulate the training procedure in Algorithm 1.
-TABLE II: Classification accuracy (%) on the CIFAR-100 validation set for different teacher-student network architecture.
-$\mathrm{\Delta}$ denotes the improvement in performance compared to vanilla KD.
-Results are averaged over five trials.
-Methods Teacher ResNet32$\times$4 WRN-40-2 VGG13 ResNet50 ResNet32$\times$4
-79.42 75.61 74.64 79.34 79.42
-Student ShuffleNetV1 ShuffleNetV1 MobileNetV2 MobileNetV2 ShuffleNetV2
-70.50 70.50 64.60 64.60 71.82
-Features FitNet[59] 73.59 73.73 64.14 63.16 73.54
-AT[31] 71.73 73.32 59.40 58.58 72.73
-RKD[55] 72.28 72.21 64.52 64.43 73.21
-CRD[70] 75.11 76.05 69.73 69.11 75.65
-OFD[24] 75.98 75.85 69.48 69.04 76.82
-ReviewKD[7] 77.45 77.14 70.37 69.89 77.78
-Logits Vanilla KD[25] 74.07 74.83 67.37 67.35 74.45
-TAKD[51] 74.53 75.34 67.91 68.02 74.82
-DKD[90] 76.45 76.70 69.71 70.35 77.07
-CKD 77.72 77.47 70.11 70.62 78.18
-$\mathrm{\Delta}$ +3.65 +2.64 +2.74 +3.27 +3.73
-IV Experiments
-IV-A Image Classification
-IV-A1 Datasets
-We first conduct experiments on CIFAR-100 [32] and ImageNet-1K [13].
-The CIFAR-100 dataset contains 100 classes, each with 600 images.
-The training and validation sets comprise 50K and 10K images.
-The ImageNet-1K dataset consists of 1000 classes.
-The training set has about 1.28 million images, and the validation set contains 50k images.
-TABLE III: Top-1 and Top-5 accuracy (%) on the ImageNet-1K validation set.
-ResNet-34 serves as the teacher, while ResNet-18 acts as the student.
+Method Homogeneous Architectures Heterogeneous Architectures
+Teacher RN56 RN110 RN32$\times$4 W-40-2 W-40-2 VGG13 RN32$\times$4 W-40-2 VGG13 RN50 RN32$\times$4
+Acc 72.34 74.31 79.42 75.61 75.61 74.64 79.42 75.61 74.64 79.34 79.42
+Student RN20 RN32 RN8$\times$4 W-16-2 W-40-1 VGG8 SN-V1 SN-V1 MN-V2 MN-V2 SN-V2
+Acc 69.06 71.14 72.50 73.26 71.98 70.36 70.50 70.50 64.60 64.60 71.82
+Feature-based Distillation
+FitNet[61] 69.21 71 06 73.50 73.58 72.24 71.02 73.59 73.73 64.14 63.16 73.54
+AT[31] 70.55 72.31 73.44 74.08 72.77 71.43 71.73 73.32 59.40 58.58 72.73
+RKD[56] 69.61 71.82 71.90 73.35 72.22 71.48 72.28 72.21 64.52 64.43 73.21
+CRD[72] 71.16 73.48 75.51 75.48 74.14 73.94 75.11 76.05 69.73 69.11 75.65
+OFD[24] 70.98 73.23 74.95 75.24 74.33 73.95 75.98 75.85 69.48 69.04 76.82
+ReviewKD[7] 71.89 73.89 75.63 76.12 75.09 74.84 77.45 77.14 70.37 69.89 77.78
+Logits-based Distillation
+KD[25] 70.66 73.08 73.33 74.92 73.54 72.98 74.07 74.83 67.37 67.35 74.45
+DML[91] 69.52 72.03 72.12 73.58 72.68 71.79 72.89 72.76 65.63 65.71 73.45
+TAKD[52] 70.83 73.37 73.81 75.12 73.78 73.23 74.53 75.34 67.91 68.02 74.82
+DKD[92] 71.97 74.11 76.32 76.24 74.81 74.68 76.45 76.70 69.71 70.35 77.07
+CKD 72.12 74.48 76.76 76.28 75.14 74.98 77.72 77.47 70.11 70.62 78.18
+$\mathrm{\Delta}$ +1.46 +1.40 +3.43 +1.36 +1.60 +2.00 +3.65 +2.64 +2.74 +3.27 +3.73
+TABLE II: Top-1 and Top-5 accuracy (%) on the ImageNet-1K validation set.(a) ResNet34 serves as the teacher, while ResNet18 acts as the student.
+(b) ResNet50 serves as the teacher, while MobileNetV1 acts as the student.
 $\mathrm{\Delta}$ denotes the performance improvement compared to vanilla KD.
 All results are the average over three trials.
-Baselines Features Logits
-Teacher Student AT[85] OFD[24] CRD[70] ReviewKD[7] Vanilla KD[25] DKD[90] CKD $\mathrm{\Delta}$
-Top-1 73.31 69.75 70.69 70.81 71.17 71.61 70.66 71.70 72.24 +1.58
+Baselines Feature-based Distillation Logits-based Distillation
+Setting Metric Teacher Student AT[87] OFD[24] CRD[72] ReviewKD[7] KD[25] DKD[92] CKD $\mathrm{\Delta}$
+(a) Top-1 73.31 69.75 70.69 70.81 71.17 71.61 70.66 71.70 72.24 +1.58
 Top-5 91.42 89.07 90.01 89.98 90.13 90.51 89.88 90.41 90.81 +0.93
-TABLE IV: Top-1 and Top-5 accuracy (%) on the ImageNet-1K validation set.
-ResNet-34 serves as the teacher, while ResNet-18 acts as the student.
-$\mathrm{\Delta}$ denotes the improved performance compared to vanilla KD.
-All results are averaged over three trials.
-Baselines Features Logits
-Teacher Student AT[85] OFD[24] CRD[70] ReviewKD[7] Vanilla KD[25] DKD[90] CKD $\mathrm{\Delta}$
-Top-1 76.16 68.87 69.56 71.25 71.37 72.56 68.58 72.05 72.97 +4.39
+(b) Top-1 76.16 68.87 69.56 71.25 71.37 72.56 68.58 72.05 72.97 +4.39
 Top-5 92.86 88.76 89.33 90.34 90.41 91.00 88.98 91.05 91.36 +2.38
-TABLE V: Object detection performance on the MS-COCO dataset.
+TABLE III: Top-1 and Top-5 accuracy (%) on the Places365 validation set.(a) ResNet34 serves as the teacher, while ResNet18 acts as the student.
+(b) ResNet50 serves as the teacher, while MobileNetV1 acts as the student.
+$\mathrm{\Delta}$ denotes the performance improvement compared to vanilla KD.
+All results are the average over three trials.
+Baselines Feature-based Distillation Logits-based Distillation
+Setting Metric Teacher Student AT[87] OFD[24] CRD[72] ReviewKD[7] KD[25] DKD[92] CKD $\mathrm{\Delta}$
+(a) Top-1 54.57 54.07 54.62 55.22 54.92 55.33 55.30 55.43 55.47 +0.17
+Top-5 85.05 84.44 84.90 85.25 85.15 85.42 85.49 85.35 85.38 -0.11
+(b) Top-1 55.51 51.74 52.05 55.19 54.71 54.83 53.07 55.25 55.51 +2.47
+Top-5 85.51 82.29 82.67 85.16 84.86 85.13 83.48 85.50 85.41 +1.93
+Given the logit matrix $\boldsymbol{T}\in\mathbb{R}^{c\times n}$, and assume that there are $c$ classes, each class containing $m$ samples.
+The negative logit space $\boldsymbol{T}_{/i}\in\mathbb{R}^{c\times(n-1)}$ for the $i$-th sample satisfies,
+$\text{rank}\left(\boldsymbol{T}_{/i}\right)\leq\min{\left(c,n-1\right)}=c,\;\textit{ s.t. }c\ll n.$ (22)
+This rank constraint implies that the effective degrees of freedom in the negative space scale with the class count $c$ rather than the sample count $n$.
+Consequently, the intrinsic dimensionality of negative pairs is governed by category correlations, reducing the required sampling complexity from $\mathcal{O}(n)$ to $\mathcal{O}(c)$.
+Furthermore, compared to class-wise contrastive learning approaches, our method achieves efficient training through well-designed triples.
+TABLE IV: Object detection performance on the MS-COCO dataset.
 We employ the two-stage Faster RCNN with FPN as the detector.
@@
 $\mathrm{\Delta}$ represents the improved performance compared to ReviewKD.
-Methods ResNet101 & ResNet18 ResNet101 & ResNet50 ResNet50 & MobileNetV2
+     Methods        RN101 & RN18        RN101 & RN50        RN50 & MN-V2
 AP AP50 AP75 AP AP50 AP75 AP AP50 AP75
@@
      Student 33.26 53.61 35.26 37.93 58.84 41.05 29.47 48.87 30.90
-Vanilla KD [25] 33.97 54.66 36.62 38.35 59.41 41.71 30.13 50.28 31.35
+     KD [25] 33.97 54.66 36.62 38.35 59.41 41.71 30.13 50.28 31.35
      FitNet[61] 34.13 54.16 36.71 38.76 59.62 41.80 30.20 49.80 31.69
@@
      $\mathrm{\Delta}$ +0.90 +1.19 +6.71 +1.09 +0.96 +1.23 +1.60 +2.20 +1.75
+TABLE V: Instance segmentation performance on the MS-COCO dataset.
+We take Mask R-CNN with FPN as the detector.
+Teacher-student pairs are ResNet101 & ResNet-18, ResNet101 & ResNet50, and ResNet50 & MobileNetV2, respectively.
+$\mathrm{\Delta}$ represents the improved performance compared to ReviewKD.
+     Method AP AP50 AP75 APS APM APL
+     RN101 38.63 60.45 41.28 19.48 41.33 55.29
+     RN18 31.25 51.07 33.10 14.18 32.80 45.53
+     ReviewKD [7] 33.62 53.91 35.96 15.03 35.31 50.30
+     CKD+ReviewKD 34.59 55.01 36.82 15.68 36.35 51.45
+     $\mathrm{\Delta}$ +0.97 +1.10 +0.86 +0.65 +1.04 +1.15
+     RN101 38.63 60.45 41.28 19.48 41.33 55.29
+     RN50 35.24 56.32 37.49 15.03 35.31 50.30
+     ReviewKD [7] 36.98 58.13 39.60 17.54 39.57 53.19
+     CKD+ReviewKD 37.89 59.45 40.61 18.40 40.62 54.91
+     $\mathrm{\Delta}$ +0.91 +1.32 +1.01 +0.86 +1.05 +1.72
+     RN50 35.24 56.32 37.49 15.03 35.31 50.30
+     MN-V2 28.37 47.19 29.95 12.09 29.01 41.70
+     ReviewKD [7] 31.56 50.70 33.44 12.76 32.44 47.39
+     CKD+ReviewKD 32.80 52.64 34.96 14.67 34.38 48.84
+     $\mathrm{\Delta}$ +1.24 +1.94 +1.52 +1.91 +1.94 +1.45
+For classic methods with $n$ samples, a total of $\mathcal{O}(nm)$ triples should be selected for each sample,
+$\binom{m-1}{1}\binom{n-m}{1}=nm-m^{2}-n+m\propto\mathcal{O}(nm)$ (23)
+Eq.
+23 shows that for each sample, these methods randomly choose one sample from the same class, including $m-1$ selections, and one sample from a different class, yielding a total of $\mathcal{O}(nm)$ triples.
+In contrast, CRD only needs to select one sample from different classes, as the positive pair is derived from the same sample,
+$\binom{1}{1}\binom{n-m}{1}=n-m\propto\mathcal{O}(n).$ (24)
+However, the negative space $\mathcal{O}(n)$ is extensive, requiring a memory bank of size $K$ to store historical features.
+There are $K$ triples available for training each sample.
+In contrast, the proposed method requires sampling only $\mathcal{O}(c)$ triples,
+$\binom{1}{1}\binom{c}{1}=c\propto\mathcal{O}(1),$ (25)
+where $c\ll n$ We randomly shuffle data samples and select $c$ non-overlapping samples sequentially.
+Therefore, our sample-wise formulation exhibits superior training efficiency compared to the class-wise contrastive formulation.
+Finally, we show the training procedure in Algorithm 1.
+IV Experiments
+IV-A Image Classification
+IV-A1 Datasets
+We conduct experiments on three benchmark datasets, including CIFAR-100 [32], ImageNet-1K [13] and Places365 [93].
+The CIFAR-100 dataset contains 100 fine-grained classes, with each class consisting of 600 images.
+The dataset is divided into 50,000 training images and 10,000 validation images.
+The ImageNet-1K dataset, a large-scale classification dataset, consists of 1000 classes, with approximately 1.28 million training images and 50,000 validation images.
+The Places365 dataset [93] is a large-scale scene recognition dataset containing 365 scene categories.
+This dataset includes 1.8 million training images and 36,500 validation images.
 IV-A2 Experimental Setup
-We evaluate eleven isomorphic or heterogeneous architectures on CIFAR-100 and compare with vanilla KD [25], AT [85], FitNet [59], RKD [55], CRD [70], OFD [24], ReviewKD [7], TAKD[51], and DKD [90].
-For ImageNet-1K, we evaluate representatives, including ResNet34 & ResNet18, and ResNet50 & MobileNetV2.
-We provide comparisons with vanilla KD [25], AT [85], OFD [24], CRD [70], ReviewKD [7], and DKD [90].
-Top-1 and top-5 recognition rates are reported.
+We evaluate eleven isomorphic or heterogeneous architectures on CIFAR-100 and comparison methods include vanilla KD [25], AT [87], FitNet [61], RKD [56], CRD [72], OFD [24], ReviewKD [7], DML[91], TAKD[52], and DKD [92].
+For the large-scale ImageNet-1K and Places365 datasets, we focus on representative architecture pairs: ResNet34 & ResNet18 and ResNet50 & MobileNetV1.
+We compares against vanilla KD [25], AT [87], OFD [24], CRD [72], ReviewKD [7], and DKD [92], reporting both top-1 and top-5 recognition rates across all experiments.
+IV-A3 Implementation Details
+All experiments are conducted on NVIDIA RTX 3090 GPUs with dataset-specific settings.
+For CIFAR-100, we set $\alpha$ to 100 and employ a cosine annealing learning rate scheduler.
+For homogeneous and heterogeneous networks, we set batch sizes to 64 and 32, with an initial learning rate of 0.05 and 0.01.
+For ImageNet-1K, we use cosine annealing with an initial learning rate of 0.2.
+We set the parameter $\alpha$ to 10 and the batch size to 512.
+For Places365, we set the parameter $\alpha$ to 10 and the batch size to 256.
 Figure 7: Comparison of temperatures on CIFAR-100. Experiments are conducted on (1) ResNet8$\times$4 and ResNet32$\times$4, (2) ShuffleNetV1 and ResNet32$\times$4.
-Figure 8: Comparison of batch sizes on CIFAR-100. Experiments are conducted on (1) ResNet8$\times$4 and ResNet32$\times$4, (2) ShuffleNetV1 and ResNet32$\times$4.
-IV-A3 Implementation Details
-All experiments are conducted on NVIDIA RTX 3090.
-For CIFAR-100, we set $\alpha$ to 100 and employ cosine annealing for learning rates.
-For homogeneous and heterogeneous networks, we set batch sizes to 64 and 32, with an initial learning rate of 0.05 and 0.01, respectively.
-For ImageNet-1K, we also use cosine annealing, with an initial learning rate of 0.2.
-We set $\alpha$ to 10 and a batch size to 512.
-Figure 9: Comparison of training time per batch and accuracy on CIFAR-100.
-The teacher model is ResNet32$\times$4 and the student model is ResNet8$\times$4.
-Additional parameters are denoted after $+$, which are added based on the student’s foundation.
 TABLE VI: Ablation studies of contrastive formulation on the CIFAR-100 dataset.
@@
 SupCon and CRD are class-wise methods, while our CKD is a sample-wise approach.
-Methods Homogeneous architectures Heterogeneous architectures
-Teacher ResNet56 WRN-40-2 WRN-40-2 VGG13 ResNet50 ResNet32$\times$4
+Method Homogeneous Heterogeneous
+Teacher RN56 W-40-2 W-40-2 VGG13 RN50 RN32$\times$4
 72.34 75.61 75.61 74.64 79.34 79.42
-Student ResNet20 WRN-16-2 WRN-40-1 MobileNetV2 MobileNetV2 ShuffleNetV2
+Student RN20 W-16-2 W-40-1 MN-V2 MN-V2 SN-V2
 69.06 73.26 71.98 64.60 64.60 71.82
@@
 CKD 72.12 76.28 75.14 70.11 70.62 78.18
+Figure 8: Comparison of batch sizes on CIFAR-100. Experiments are conducted on (1) ResNet8$\times$4 and ResNet32$\times$4, (2) ShuffleNetV1 and ResNet32$\times$4.
 TABLE VII: Ablation studies of triple design on the CIFAR-100 dataset.
 Experiments are conducted on both homogeneous and heterogeneous teacher-student pairs.
-Due to the fixed teacher, triples $\left(\mathbf{t}_{i},\mathbf{s}_{i},\mathbf{t}_{j}\right)$ is meaningless and excluded from this table.
-Triples Homogeneous architectures Heterogeneous architectures
-Teacher ResNet56 WRN-40-2 WRN-40-2 VGG13 ResNet50 ResNet32$\times$4
+Due to the fixed teacher, triples $\left(\mathbf{t}_{i},\mathbf{s}_{i},\mathbf{t}_{j}\right)$ are meaningless and excluded from this table.
+Triple Homogeneous Heterogeneous
+Teacher RN56 W-40-2 W-40-2 VGG13 RN50 RN32$\times$4
 72.34 75.61 75.61 74.64 79.34 79.42
-Student ResNet20 WRN-16-2 WRN-40-1 MobileNetV2 MobileNetV2 ShuffleNetV2
+Student RN20 W-16-2 W-40-1 MN-V2 MN-V2 SN-V2
 69.06 73.26 71.98 64.60 64.60 71.82
-CKD$\left(\mathbf{s}_{i},\mathbf{t}_{i},\mathbf{s}_{j}\right)$ 71.76 75.31 74.40 70.04 68.93 78.02
-CKD$\left(\mathbf{s}_{i},\mathbf{t}_{i},\mathbf{t}_{j}\right)$ 71.69 75.88 74.43 70.09 69.68 77.06
-CKD$\left(\mathbf{t}_{i},\mathbf{s}_{i},\mathbf{s}_{j}\right)$ 72.12 76.28 75.14 70.11 70.62 78.18
+$\left({s}_{i},{t}_{i},{s}_{j}\right)$ 71.76 75.31 74.40 70.04 68.93 78.02
+$\left({s}_{i},{t}_{i},{t}_{j}\right)$ 71.69 75.88 74.43 70.09 69.68 77.06
+$\boldsymbol{\left({t}_{i},{s}_{i},{s}_{j}\right)}$ 72.12 76.28 75.14 70.11 70.62 78.18
 IV-A4 Performance on CIFAR-100
-We conduct classification experiments by using eleven combinations of isomorphic or heterogeneous teacher and student network architectures, including ResNet [22], WideResNet [73], VGG [63], ShuffleNet [88, 47], and MobileNet [60].
-Table I presents classification performance by using teacher and student architectures of the same style.
-We can observe that our method achieves competitive performance against feature-based and logit-based methods.
-Specifically, the proposed CKD improves vanilla KD accuracy by 1.36%, with a maximum gain of 3.43%.
-Moreover, CKD also outperforms CRD on the CIFAR-100 dataset, with an improvement of almost 1.00%.
-Table II shows the classification performance of teacher-student networks with different styles.
-We see that our CKD method consistently yields improvements across teacher-student network pairs compared to vanilla KD.
-Specifically, for different architectures, CKD achieves an average improvement of 3.21% over vanilla KD, with an impressive maximum gain of 3.73%.
+We conduct classification experiments by using eleven combinations of isomorphic or heterogeneous teacher-student networks, including ResNet [22], WideResNet [75], VGG [65], ShuffleNet [90, 47], and MobileNet [62].
+Table I shows classification performance using both homogeneous and heterogeneous teacher-student architectures.
+Our method achieves competitive performance against feature-based and logit-based methods.
+The proposed CKD achieves an average improvement of 2.20% over vanilla KD across all 11 settings, with a maximum gain of 3.73%.
+Our method surpasses state-of-the-art feature distillation approaches, with an average accuracy improvement of 1.00% against CKD.
+Moreover, CKD shows stronger improvements in heterogeneous settings compared to homogeneous configurations.
+The superior performance of sample-wise contrastive learning over class-wise alternatives is evident in heterogeneous scenarios.
+The sample-wise contrastive mechanism appears particularly effective when dealing with heterogeneous architectures.
+Unlike feature-based methods requiring explicit layer alignment, CKD creates architecture-agnostic embeddings with logit alignment, better bridging the representation gap between different networks.
 IV-A5 Performance on ImageNet-1K
-To further verify the ability of CKD, we perform experiments on the ImageNet-1K dataset.
-Table III and Table IV report experimental results on the ImageNet-1K dataset.
-We can see that our method achieves superior performance with both teacher-student architectures.
-Moreover, the performance of CKD surpasses state-of-the-art results, providing further evidence of the effectiveness of our approach.
-Specifically, the proposed CKD achieves accuracy improvements of 1.58% and 4.39% over vanilla KD with the same and different teacher-student architectures.
-IV-B Object Detection
+To further verify the effectiveness of CKD on a large-scale dataset, we perform experiments on ImageNet-1K.
+As listed in Table II, our method achieves superior performance with both teacher-student architectures, providing further evidence of the effectiveness on large-scale datasets.
+Specifically, when distilling between homogeneous networks, CKD achieves 72.24% Top-1 accuracy, representing a 1.58% improvement over vanilla KD.
+More remarkably, in the cross-architecture setting, CKD attains 72.97% Top-1 accuracy with a substantial 4.39% gain.
+These results not only validate its compatibility with diverse network structures but also establish new state-of-the-art performance on large-scale knowledge distillation.
+IV-A6 Performance on Places365
+To evaluate our method on large-scale scene recognition, we conduct experiments on the Places365 validation set.
+As shown in Table III, our method achieves the SOTA Top-1 accuracy and competitive Top-5 accuracy, demonstrating its effectiveness in complex scene understanding tasks.
+Notably, our method outperforms CRD in all settings, highlighting the superiority of sample-wise contrastive learning over class-wise approaches.
+These results validate the robustness and scalability of our method in large-scale knowledge distillation tasks.
+IV-B Object Detection and Instance Segmentation
 IV-B1 Datasets
-We conduct experiments on the MS-COCO dataset [43].
-MS-COCO is a popular dataset featuring 80 categories, divided into two subsets, including a training set of 118,000 images and a validation set of 5,000 images.
+We evaluate our method on the MS-COCO dataset [43], which contains 118,000 training and 5,000 validation images across 80 object categories.
+Figure 9: Comparison of training time per batch and accuracy on CIFAR-100.
+The teacher model is ResNet32$\times$4, and the student model is ResNet8$\times$4.
+Additional parameters are denoted after $+$ and are added based on the student’s foundation.
 IV-B2 Experimental Setup
-We analyze ResNet-101 & ResNet-18, ResNet-101 & ResNet-50, and ResNet-50 & MobileNetV2 teacher-student pairs. Moreover, we evaluate the MS-COCO validation set with the baseline provided by Detectron2.
+We analyze ResNet-101 & ResNet-18, ResNet-101 & ResNet-50, and ResNet-50 & MobileNetV2 teacher-student pairs.
+Following standard evaluation protocols, we implement the detection framework using Detectron2.
 IV-B3 Implementation Details
-Experiments are performed on NVIDIA RTX 3090.
-We utilize Faster R-CNN with FPN as the feature extractor and implement CKD on the R-CNN head.
+All experiments are conducted on NVIDIA RTX 3090 GPUs.
+For object detection, we employ Faster R-CNN with FPN as the feature extractor and implement CKD on the R-CNN head.
+For instance segmentation, we utilize Mask R-CNN as the baseline architecture.
 We set $\alpha$ to 0.1 and the temperature $\tau$ to 1.0.
-Schedulers and task-specific loss settings are following Detectron2.
-IV-B4 Performance on MS-COCO
-Table V presents the object detection results on the MS-COCO dataset under three teacher-student settings.
-We can see that the proposed approach obtains comparable performance against logit-based approaches but inferior performance compared to feature-based approaches, particularly for ReviewKD.
-Object detection relies on spatial representations to locate objects.
+All training protocols, including learning rate schedulers and loss functions, strictly follow Detectron2.
+IV-B4 Object Detection Performance on MS-COCO
+Table IV presents the object detection results on the MS-COCO dataset under three teacher-student settings.
+We can see that the proposed CKD achieves comparable performance against logit-based approaches but inferior performance compared to feature-based approaches, particularly for ReviewKD.
+This can be attributed to the fact that object detection heavily relies on spatial representations to locate objects.
 However, logits lack detailed spatial information [36, 78].
-ReviewKD [7] conducts multi-level feature distillation, resulting in better performance.
-We combine ReviewKD with logit-based approaches to utilize spatial information.
-Notably, the combined approaches, including CKD+ReviewKD and DKD+ReviewKD, significantly improve object detection results compared to previous counterparts.
+ReviewKD [7] conducts multi-level feature distillation, leading to superior performance.
+To leverage spatial information, we combine ReviewKD with logit-based approaches.
+Notably, the combined approaches, including CKD+ReviewKD and DKD+ReviewKD, significantly improve detection results compared to their standalone counterparts.
+IV-B5 Instance Segmentation Performance on MS-COCO
+Table V presents instance segmentation results on the MS-COCO dataset across different frameworks.
+Notably, existing knowledge distillation methods specifically designed for instance segmentation remain scarce, which limits comparisons primarily to ReviewKD and the proposed CKD+ReviewKD.
+CKD+ReviewKD achieves performance gains over ReviewKD, with absolute accuracy increases of +0.97, +0.91, and +1.24 points in different teacher-student pairs, respectively.
+The most significant improvement emerges in the RN50-MobileNetV2 setting, particularly for small objects.
+The synergy between logit-based distillation CKD and feature alignment ReviewKD addresses both global semantic preservation and local spatial detail transfer.
+This dual mechanism explains the balanced improvements across all object scales [36].
+Experiments demonstrate robustness across different network gaps from RN101-RN50 to RN50-MN-V2, suggesting wide applicability in practical scenarios where compact models are essential.
 IV-C Ablation Study
-IV-C1 Effects of Temperature $\tau$
+IV-C1 Effect of Temperature $\tau$
 In the training phase, a high $\tau$ yields a smooth distribution across classes, making distillation more challenging.
-But, a low $\tau$ leads to a sharp distribution.
-To investigate the effect of $\tau$ on CKD, we conduct ablation studies by using ResNet32×$\times 4$ ResNet8×$\times 4$ and ResNet32×$\times 4$ ShuffleNetV1 pairs.
+But a low $\tau$ results in a sharp distribution.
+To investigate the effect of $\tau$, we conduct ablation studies by using ResNet32×$\times 4$ ResNet8×$\times 4$ and ResNet32×$\times 4$ ShuffleNetV1 pairs.
 Typically, $\tau$ is set to a low value within a range of 0.07 to 4.0, while other parameters are kept constant.
@@
 WRN-16-2 is used as a student.
-In the left-hand section of the table, WRN series networks serve as teachers, while in the right-hand section, different series networks are utilized as teachers.
-Methods WRN Series Different Series
-Teacher WRN-40-2 WRN-16-4 WRN-28-4 VGG13 WRN-16-4 ResNet50
+In the left-hand section, WRN series networks serve as teachers, while in the right-hand section, different series networks are utilized as teachers.
+Method WRN Series Different Series
+Teacher W-40-2 W-16-4 W-28-4 VGG13 W-16-4 RN50
 75.61 77.21 79.12 74.64 77.21 79.34
 Studnet 73.26 73.26 73.26 73.26 73.26 73.26
-vanilla KD[25] 74.92 76.00 75.14 74.30 76.00 74.62
+KD[25] 74.92 76.00 75.14 74.30 76.00 74.62
 CKD 76.28 76.22 76.42 75.14 76.22 76.45
 Fig.
-7 presents Top-1 classification accuracies for different temperatures $\tau$.
-We can observe that our method performs best when temperature $\tau$ is around 1.0 and performs poorly when $\tau$ is low or high.
-This suggests that our method is better suited for aligning ${t}_{i}$ and ${s}_{i}$ when the value of $\tau$ is near 1.0.
+7 illustrates the temperature sensitivity analysis for distillation.
+Both settings exhibit a clear performance parabola with peak Top-1 accuracy 79.0% and 76.3%, respectively, at optimal temperature $\tau=1.0$.
+Our method achieves optimal performance when temperature $\tau$ is approximately 1.0 and performs poorly when $\tau$ is low or high.
+This suggests that proper temperature scaling balances the sharpness of knowledge transfer.
 From Theorem 1, we conclude that the gradient $\nabla_{\boldsymbol{s}_{*}}\mathcal{L}_{i}$ is proportional to $g_{i}$.
-When $\tau<1$, the difference between $e^{f\left(\boldsymbol{t}_{i},\boldsymbol{s}_{i}\right)/\tau}$ and $e^{f\left(\boldsymbol{t}_{i},\boldsymbol{s}_{j}\right)/\tau}$ is magnified exponentially.
+When $\tau<1$, the difference between $e^{f\left(\boldsymbol{t}_{i},\boldsymbol{s}_{i}\right)/\tau}$ and $e^{f\left(\boldsymbol{t}_{i},\boldsymbol{s}_{j}\right)/\tau}$ is exponentially magnified.
 In the limit, we obtain $\lim_{\tau\to 0}g_{i}=0$.
-As a result, $g_{i}$ approaches zero, causing a rapid decline in performance.
-When $\tau>1$, the difference is lessened and $g_{i}$ becomes more sensitive to negative samples as $\lim_{\tau\to\infty}g_{i}=\frac{B}{B+1}$.
-IV-C2 Effects of Batchsize
-To analyze the effect of batch size, we utilize two teacher-student pairs including ResNet56 & ResNet20 and ResNet32×$\times 4$ ShuffleNetV2.
-Moreover, we set the batch size to $2^{k}$, where $k\in[4,11]\cap k\in\mathbb{N}$.
+As a result, $g_{i}$ approaches zero, causing a vanishing gradient.
+When $\tau>1$, the difference is reduced and $g_{i}$ becomes more sensitive to negative samples, as $\lim_{\tau\to\infty}g_{i}=\frac{B}{B+1}$.
+Figure 10: Class Activation Map via GradCAM. The teacher networks (ResNet34) and the student networks (ResNet18) are trained on ImageNet-1K.
+IV-C2 Effect of Batchsize
+To analyze the effect of batch size, we utilize two teacher-student pairs: ResNet56 & ResNet20 and ResNet32×$\times 4$ ShuffleNetV2.
+We set the batch size $B$ to $2^{k}$, where $k\in[4,11]\cap k\in\mathbb{N}$.
 Fig.
-8 presents Top-1 classification performance with varying batch sizes.
-We can see that our method attains good performance when the batch size is 64 for ResNet56 & ResNet20 and 32 for ResNet32×$\times 4$ ShuffleNetV2.
-Larger or smaller batch sizes are not optimal.
-We also find that a large batch size potentially hampers the efficacy of our method as it affects intra-sample alignment through an increased $\beta$.
-However, a small batch size can lead to insufficient gradient magnitude.
-In fact, the optimal batch size for the sample-wise formulation is strongly correlated with the number of categories.
-It should be smaller than the number of categories to avoid neglecting intra-sample distillation and facilitate effective inter-sample distillation.
-Figure 10: Class Activation Map via GradCAM. The teacher networks (ResNet34) and the student networks (ResNet18) are trained on ImageNet-1K.
+8 presents top-1 accuray across various batch sizes.
+We can see that our method achieves maximum accuracies when the batch size is 64 for ResNet56 & ResNet20 and 32 for ResNet32×$\times 4$ ShuffleNetV2.
+Batch sizes outside these optimal values yield suboptimal results.
+Experiments reveal that a large batch size potentially hampers the efficacy of our method by overwhelming inter-sample distillation through an increased $\beta$.
+However, a small batch size suffers from insufficient negative pairs due to vanishing gradients.
+The empirical optima should satisfy the category count constraint $B<c$.
+When $B>c$, distillation focuses on inter-sample relations, degrading class-specific knowledge transfer.
+This analysis proves that batch size significantly balances intra/inter-sample knowledge transfer.
 IV-C3 Influence of Contrastive Formulation
 Fig.
-6 visualizes class-wise and sample-wise contrastive formulations.
-SupCon [28] is a class-wise approach that selects positive pairs within the same class.
-Noat that CRD [70] is a typical class-wise method with positive pairs from the same sample and negative pairs from different classes.
-In contrast, CKD chooses negative pairs from different samples but potentially from the same class.
-Table VI presents classification performance under different formulations.
-We can observe that CKD outperforms SupCon and CRD because it uses sample-wise negative samples.
-Moreover, CRD exploits a memory bank to access a large number of negative samples and achieves a higher accuracy than SupCon.
+6 visualizes differences between class-wise and sample-wise contrastive learning frameworks.
+In class-wise approaches such as SupCon [28], positive pairs are formed exclusively within the same class.
+This category also includes methods like CRD [72], which specifically constructs positive pairs from teacher-student models of the same sample while selecting negative pairs from different classes.
+By contrast, the sample-wise CKD adopts a different strategy in that it selects negative pairs from distinct samples, even when they potentially belong to the same class.
+Table VI compares classification performance across three contrastive learning frameworks.
+Notably, CRD exhibits improved accuracy over SupCon by maintaining a memory bank to access negative instances.
+However, our CKD achieves superior performance by employing sample-wise negative pairs, which is particularly evident in heterogeneous settings where it outperforms CRD by 2.53% with the SN-v2 setting.
 IV-C4 Influence of Triple Design
-In knowledge distillation, the element of triples can be sampled from teacher or student models.
-Hence, there are several kinds of triples, including $\left(\boldsymbol{s}_{i},\boldsymbol{t}_{i},\boldsymbol{s}_{j}\right)$, $\left(\boldsymbol{s}_{i},\boldsymbol{t}_{i},\boldsymbol{t}_{j}\right)$, $\left(\boldsymbol{t}_{i},\boldsymbol{s}_{i},\boldsymbol{t}_{j}\right)$ and $\left(\boldsymbol{t}_{i},\boldsymbol{s}_{i},\boldsymbol{s}_{j}\right)$.
+In knowledge distillation, triplet elements can be sampled from either the teacher (t) or student (s) models.
+Hence, there are four kinds of triples, including $\left(\boldsymbol{s}_{i},\boldsymbol{t}_{i},\boldsymbol{s}_{j}\right)$, $\left(\boldsymbol{s}_{i},\boldsymbol{t}_{i},\boldsymbol{t}_{j}\right)$, $\left(\boldsymbol{t}_{i},\boldsymbol{s}_{i},\boldsymbol{t}_{j}\right)$ and $\left(\boldsymbol{t}_{i},\boldsymbol{s}_{i},\boldsymbol{s}_{j}\right)$.
 $\left(\boldsymbol{s}_{i},\boldsymbol{t}_{i},\boldsymbol{s}_{j}\right)$ means that we align $\mathbf{s}_{i}$ and $\boldsymbol{t}_{i}$ while distancing $\boldsymbol{s}_{i}$ from $\boldsymbol{s}_{j}$.
-This triple can be implemented by modifying diagonal elements of the affinity matrix with $\boldsymbol{s}_{i}$ and $\boldsymbol{t}_{i}$.
-Due to fixed teacher logits, $\left(\boldsymbol{t}_{i},\boldsymbol{s}_{i},\boldsymbol{t}_{j}\right)$ equals intra-sample alignment.
-Table VII shows the classification performance with different triples.
-We can see that the utilization of triples $\left(\boldsymbol{t}_{i},\boldsymbol{s}_{i},\boldsymbol{s}_{j}\right)$ yields better performance than the other triples.
-Our method anchors on $\boldsymbol{t}_{i}$, pulling $\boldsymbol{s}_{i}$ closer while pushing $\boldsymbol{s}_{j}$ apart.
-This objective is clear and simple.
-However, as discussed in Section III-B, triples $\left(\boldsymbol{s}_{i},\boldsymbol{t}_{i},\boldsymbol{t}_{j}\right)$ may cause gradient conflict.
-For triples $\left(\boldsymbol{s}_{i},\boldsymbol{t}_{i},\boldsymbol{s}_{j}\right)$, this strategy mainly performs contrastive learning in the student logits space.
+This triple is implemented through diagonal modification of the affinity matrix between $\boldsymbol{s}_{i}$ and $\boldsymbol{t}_{i}$ pairs.
+Notably, $\left(\boldsymbol{t}_{i},\boldsymbol{s}_{i},\boldsymbol{t}_{j}\right)$ reduces to intra-sample alignment since teacher logits remain fixed during distillation..
+Table VII shows the classification performance for three triple settings.
+The teacher-anchored triples $\left(\boldsymbol{t}_{i},\boldsymbol{s}_{i},\boldsymbol{s}_{j}\right)$ demonstrates superior performance across all settings, achieving 1.43% higher accuracy than $\left(\boldsymbol{s}_{i},\boldsymbol{t}_{i},\boldsymbol{t}_{j}\right)$ in SN-V2 heterogeneous architecture.
+This advantage stems from its clear optimization objective by anchoring on fixed teacher features $\boldsymbol{t}_{i}$ to simultaneously attract student $\boldsymbol{s}_{i}$ and repel $\boldsymbol{s}_{j}$.
+However, as discussed in Section III-B, triples $\left(\boldsymbol{s}_{i},\boldsymbol{t}_{i},\boldsymbol{t}_{j}\right)$ risks gradient conflict due to competing objectives between moving $\boldsymbol{s}_{i}$ toward $\boldsymbol{t}_{i}$ and away from $\boldsymbol{t}_{j}$.
+For triples $\left(\boldsymbol{s}_{i},\boldsymbol{t}_{i},\boldsymbol{s}_{j}\right)$, this strategy mainly performs contrastive learning in the student logits space, yet insufficiently leverages the prior knowledge from the teacher model.
 IV-D More Analyses
 IV-D1 Training Efficiency
-To evaluate the training efficiency of our approach, we compare its accuracy and parameter count against state-of-the-art methods on the CIFAR-100 dataset.
+To evaluate the efficiency-accuracy trade-off, we compare our method against knowledge distillation baselines on CIFAR-100.
+As shown in Fig.
+9, our method exhibits high training efficiency, striking a good balance between model parameters and classification performance.
+Feature-based methods require more parameters and longer training times than logit-based methods.
+ReviewKD and CRD, in particular, take twice the training time over our method due to multi-level feature review and memory bank.
+Our method simultaneously improves accuracy and reduces training time.
+IV-D2 Effect of Teacher Capacity
+Existing studies [10, 76] have demonstrated that large teacher models may be suboptimal for distilling knowledge into smaller student models due to parameter and capacity gaps between the two architectures.
+To investigate this phenomenon, we conduct experiments to examine how teacher networks with varying model capacities affect knowledge transfer efficiency.
+Table VIII compares knowledge distillation performance on the CIFAR-100 dataset.
+We can see that our method achieves comparable performance with increasing WRN teacher capacity, whereas vanilla KD exhibits variation in performance.
+For vanilla KD, larger models lead to performance degradation, achieving the accuracies of WRN-16-4 and WRN-28-4 at 76.00 and 75.14, respectively.
+In contrast, our CKD achieves progressive accuracy improvements as teacher capacity increases from WRN-40-2 to WRN-28-4, demonstrating effective utilization of enhanced teacher representations.
+Wide Residual Networks [75] is a ResNet variant.
+Our method also obtains an accuracy on par with ResNet50 as the teacher model, while vanilla KD has the second-worst performance.
+Both methods show limited effectiveness with VGG13 architecture, suggesting architectural compatibility influences distillation efficacy.
+These results prove that our method can partially mitigate the parameter and capacity gaps of larger models.
+IV-E Visualizations
 Fig.
-9 depicts training time per batch and classification performance.
-We can observe that our method exhibits high training efficiency, striking a balance between model parameters and classification performance.
-Typically, feature-based methods require more parameters and longer training times than logit-based methods.
-ReviewKD and CRD, in particular, take twice the training time compared to our method, due to multi-level feature review and memory bank usage.
-IV-D2 Effects of Teacher Capacity
-Previous works [10, 74] have found that large teacher models may not be optimal for smaller student models due to disparities in parameters and capacity.
-Therefore, we further conduct experiments to analyze the influence of teacher networks with different capacities.
-Table VIII presents classification performance on the CIFAR-100 dataset.
-We can see that our method achieves comparable performance across different WRN models, whereas vanilla KD exhibits variation in performance.
-Wide Residual Networks [73], is a ResNet variant.
-Therefore, our method also obtains an accuracy on par with ResNet50 as the teacher model, while vanilla KD has the second-worst performance.
-For vanilla KD, larger models lead to performance degeneration, for example, the accuracies of WRN-16-4 and WRN-28-4 are 76.00 and 75.14, respectively.
-Both the proposed method and vanilla KD perform poorly with the VGG13 architecture.
-Ablation studies prove that our method can partially mitigate the parameter and capacity gaps of larger models.
-IV-E Visualizations
-GradCAM identifies image regions critical for concept prediction.
-To evaluate the consistency of CKD with teacher networks, we employ GradCAM to visualize class activation maps in Fig.
-10.
-The teacher-student pair, including ResNet34 and ResNet18, are trained on the ImageNet-1K dataset.
+10 demonstrates the semantic alignment between CKD and teacher networks through GradCAM visualizations, where the ResNet34-ResNet18 pair was trained on ImageNet-1K.
 We can observe that our method effectively handles complex scenes and multiple objects.
-For example, in the seventh column, two birds are present.
-Our method accurately replicates teacher knowledge by focusing on both birds, whereas student networks and vanilla KD focus only on the bird closer to the camera.
+In multi-object cases, CKD captures both discriminative features of birds circled in red, while the student baseline and vanilla KD focus solely on the foreground bird.
+Across diverse scenes like natural landscapes and indoor settings, CKD consistently preserves the spatial attention patterns, particularly in fine-grained regions like butterfly wing patterns.
+In the final column, CKD aligns with the teacher model in localizing the target dog, whereas the student network and vanilla KD erroneously emphasize the distractor dog.
+These visual evidence confirm that CKD maintains teacher-aligned sample-specific semantics rather than generic category features.
 V Conclusion
@@
 By converting the structure-preserving alignment into a sample-wise contrastive learning framework, our approach can be effectively and efficiently trained with well-designed positive and negative pairs.
-Experimental results demonstrate that our method achieves highly competitive performance against state-of-the-art methods on benchmark datasets.
-In the future, we will attempt to apply the proposed method to distill large models.
+Experiments demonstrate that our method achieves highly competitive performance against state-of-the-art methods on benchmarks.
+In the future, we will explore the application of this framework to large multimodal models.